NEWS

Geoff Huston Last Revised: 9 January 1989

Software Version: V5.7

April 13, 1989

Contents

1 Abstract 16 1.1 Author . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16

2 Introduction 17 2.1 Sources of Further Information . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18

3 The User Interface to NEWS 19 3.1 Invoking NEWS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 3.2 NEWS Context . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 3.3 The Structure of NEWS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 3.4 NEWS Screen Formats . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21

3.4.1 NEWSGROUP Directory Screen . . . . . . . . . . . . . . . . . . . . . . . 21 3.4.2 NEWSITEM Directory Screen . . . . . . . . . . . . . . . . . . . . . . . . 22 3.4.3 NEWS ITEM Text Display . . . . . . . . . . . . . . . . . . . . . . . . . . 23

4 ( 24 4.1 Default Keypad Definitions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 4.2 NEWS Screen Displays . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 4.3 Displaying NEWS text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 4.4 Posting Items into NEWS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 4.5 Deleting Newsitems . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 4.6 NEWS Interface to MAIL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 4.7 Extraction of Newsitems . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 4.8 Keeping Track of Newsitems . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 4.9 HELP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 4.10 EXITing NEWS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28

5 ( 29 5.1 Key Definition . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 5.2 Logical Names . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29

5.2.1 NEWS_EDIT Command Procedure . . . . . . . . . . . . . . . . . . . . . 31

6 NEWS Commands - User Reference 33

6.1 $ NEWS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 6.1.1 /SCAN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 6.1.2 /ALLSCAN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 6.1.3 /UNSEEN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 6.1.4 /SCREEN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 6.1.5 Parameters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34

6.2 ADD ENTRY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36

6.2.1 /CLASSES . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 6.3 ADD FILE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37

6.3.1 /CREGRP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 6.3.2 /DELETE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 6.3.3 /EXECUTE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 6.3.4 /DEFNEWSGROUP=newsgroup . . . . . . . . . . . . . . . . . . . . . . . 39 6.3.5 /NETFEED . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 6.3.6 /JUNK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 6.3.7 /ACCEPT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39

6.3.8 /RETRY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40

6.3.9 /MODERATOR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40

6.4 ADD MARKER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41

6.5 ADD MEMBER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42

6.5.1 /MAIL=mail-address . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42

6.5.2 /PRIVILEGES=priv-list . . . . . . . . . . . . . . . . . . . . . . . . . . . 42

6.6 ANSWER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43

6.6.1 /HEADERS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43

6.7 ATTACH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44

6.8 BACK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45

6.8.1 /HEADER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45

6.8.2 /EDITOR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45

6.8.3 /ROT13 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45

6.9 BACK REPLY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46

6.10 BACK NOTE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47

6.11 CANCEL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48

6.12 CLEAR KILL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49

6.13 CLEAR MARK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50

6.14 CLOSE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51

6.15 CREATE NEWSGROUPS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52

6.15.1 /HOLD=n . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52

6.15.2 /ITEMHOLD=n . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52

6.15.3 /MODERATOR=address . . . . . . . . . . . . . . . . . . . . . . . . . . . 52

6.15.4 /LOCAL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53

6.15.5 /RESTRICT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 6.15.6 /SERVER=node . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53

6.15.7 /PROTOCOL=proto . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54

6.15.8 /KEEPREQUEST . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54

6.15.9 /FILEHOLD=n . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54

6.15.10 /INQUIRE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54

6.15.11 /SINCE=date-timne . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55

6.16 DEFINE/KEY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56

6.16.1 /ECHO . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56

6.16.2 /PROTECT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57

6.16.3 /IF_STATE=(state-name,...) . . . . . . . . . . . . . . . . . . . . . . . . . 57

6.16.4 /LOCK_STATE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57

6.16.5 /SET_STATE=state-name . . . . . . . . . . . . . . . . . . . . . . . . . . 57

6.16.6 /TERMINATE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57

6.17 DELETE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58

6.18 DELETE ENTRY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59

6.18.1 /CLASS=class-name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59

6.19 DELETE NEWSGROUPS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60

6.19.1 /CONFIRM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60

6.20 DELETE NEWSITEMS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61

6.20.1 /CONFIRM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61

6.20.2 /IDENTIFIER=”message-id” . . . . . . . . . . . . . . . . . . . . . . . . . 61

6.21 DELETE MARKER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62

6.22 DELETE MEMBER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63

6.23 DEREGISTER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64

6.23.1 /ALL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64

6.24 DIRECTORY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 6.24.1 /CLASS=class-name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65

6.24.2 /ALL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65

6.24.3 /NEW . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65

6.24.4 /REGISTER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65

6.24.5 /SINCE=date . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65

6.24.6 /ITEMS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66

6.24.7 /NEWSGROUPS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66

6.24.8 /RESET . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66

6.25 DOWN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67

6.26 EDIT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68

6.27 EXIT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69

6.28 EXTRACT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70

6.28.1 /ALL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70

6.28.2 /APPEND . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70

6.28.3 /AUTHOR=mailaddress . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70

6.28.4 /BEFORE=date-time . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70

6.28.5 /FROM=mailaddress . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70

6.28.6 /HEADER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70

6.28.7 /MARKER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71

6.28.8 /SEEN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71

6.28.9 /SINCE=date-time . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71

6.28.10 /SUBJECT=”string” . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71

6.28.11 /TITLE=”string” . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71

6.28.12 /UNREAD . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71

6.28.13 /UNSEEN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72

6.29 FOLLOWUP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73

6.29.1 /HEADERS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73

6.30 FORWARD . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74

6.30.1 /TO=user . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74

6.30.2 /SUBJECT=”subject” . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74

6.30.3 /NOEDIT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74

6.30.4 /SELF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75

6.30.5 /HEADERS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75

6.31 HELP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76

6.32 KILL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77

6.32.1 /FROM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77

6.32.2 /SUBJECT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77

6.32.3 /HEADER=”text” . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77

6.33 MAIL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79

6.34 MARK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80

6.35 MODIFY ENTRY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81

6.35.1 /CLASSES=(class-list) . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81

6.36 MODIFY KILL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82

6.37 MODIFY MEMBER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83

6.37.1 /MAIL=”mail-address” . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83

6.37.2 /NAME=new-username . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83

6.37.3 /PRIVILEGES=(new-priv-list) . . . . . . . . . . . . . . . . . . . . . . . . 83

6.38 NEWSGROUPS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84

6.39 NOSCREEN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85

6.40 OPEN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86

6.40.1 /NEW . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86

6.40.2 /MARK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 6.41 POST . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87

6.41.1 /EDIT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87

6.41.2 /HEADERS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87

6.41.3 /NEWSGROUPS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88

6.41.4 /SELF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88

6.41.5 /SUBJECT=subj . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88

6.42 PRINT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89

6.42.1 /WAIT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89

6.42.2 /DIRECTORY=option [mode] . . . . . . . . . . . . . . . . . . . . . . . . 89

6.42.3 /NEWSGROUPS=(group-list) [directory] . . . . . . . . . . . . . . . . . . 90

6.42.4 /ALL [item] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90

6.42.5 /AUTHOR=mailaddress [item] . . . . . . . . . . . . . . . . . . . . . . . . 90

6.42.6 /BEFORE=date-time [item] . . . . . . . . . . . . . . . . . . . . . . . . . 90

6.42.7 /CONFIRM [item] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90

6.42.8 /FROM=mailaddress [item] . . . . . . . . . . . . . . . . . . . . . . . . . . 90

6.42.9 /HEADER [item] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90

6.42.10 /MARKER [item] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91

6.42.11 /SEEN [item] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91

6.42.12 /SINCE=date-time [item] . . . . . . . . . . . . . . . . . . . . . . . . . . . 91

6.42.13 /SUBJECT=”string” [item] . . . . . . . . . . . . . . . . . . . . . . . . . . 91

6.42.14 /TITLE=”string” [item] . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91

6.42.15 /UNREAD [item] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91

6.42.16 /UNSEEN [item] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91

6.42.17 /AFTER=date-time [print] . . . . . . . . . . . . . . . . . . . . . . . . . . 92

6.42.18 /BURST=ALL—ONE [print] . . . . . . . . . . . . . . . . . . . . . . . . . 92

6.42.19 /CHARACTERISTICS=char-value [print] . . . . . . . . . . . . . . . . . . 92 6.42.20 /COPIES=n [print] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92

6.42.21 /FEED [print] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92

6.42.22 /FLAG=ALL—ONE [print] . . . . . . . . . . . . . . . . . . . . . . . . . . 92

6.42.23 /FORM=form [print] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92

6.42.24 /HOLD [print] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92

6.42.25 /JOBCOUNT=n [print] . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92

6.42.26 /LOWERCASE [print] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93

6.42.27 /NAME=job-name [print] . . . . . . . . . . . . . . . . . . . . . . . . . . . 93

6.42.28 /NOTE=”job-note” [print] . . . . . . . . . . . . . . . . . . . . . . . . . . 93

6.42.29 /NOTIFY [print] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93

6.42.30 /OPERATOR=”operator request” [print] . . . . . . . . . . . . . . . . . . 93

6.42.31 /PASSALL [print] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93

6.42.32 /PARAMETERS=(param-list) [print] . . . . . . . . . . . . . . . . . . . . 93

6.42.33 /PRIORITY=priority [print] . . . . . . . . . . . . . . . . . . . . . . . . . 93

6.42.34 /QUEUE=queue-name [print] . . . . . . . . . . . . . . . . . . . . . . . . . 93

6.42.35 /SETUP=setup-modules [print] . . . . . . . . . . . . . . . . . . . . . . . . 94

6.42.36 /TRAILER=ALL—ONE [print] . . . . . . . . . . . . . . . . . . . . . . . 94

6.43 QUIT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95

6.44 READ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96

6.44.1 READ newsitem_number . . . . . . . . . . . . . . . . . . . . . . . . . . . 96

6.44.2 /HEADER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96

6.44.3 /NEW . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96

6.44.4 /EDITOR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96

6.44.5 /TPU . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96

6.44.6 /NEXT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97

6.44.7 /LAST . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 6.44.8 /BACK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97

6.44.9 /PARENT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97

6.44.10 /IDENTIFIER=”id” . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97

6.44.11 /ROT13 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97

6.44.12 /MARKER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98

6.44.13 /FOLLOWUP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98

6.44.14 /PREV . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98

6.44.15 /TOPIC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98

6.45 REFRESH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99

6.46 REGISTER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100

6.46.1 /CONFIRM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100

6.46.2 /ALL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100

6.46.3 /BEFORE=date-time . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100

6.46.4 /LOCAL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100

6.46.5 /NETGROUPS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100

6.46.6 /NEW . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101

6.46.7 /PRIORITY=n . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101

6.46.8 /RECORD . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101

6.46.9 /SINCE=date-time . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101

6.47 REPLY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102

6.47.1 /HEADERS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102

6.47.2 /SUBJECT=”mail-subject” . . . . . . . . . . . . . . . . . . . . . . . . . . 102

6.47.3 /SELF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102

6.48 SAVE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103

6.48.1 /ALL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103

6.48.2 /APPEND . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 6.48.3 /AUTHOR=mailaddress . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103

6.48.4 /BEFORE=date-time . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103

6.48.5 /FROM=mailaddress . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103

6.48.6 /HEADER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103

6.48.7 /MARKER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104

6.48.8 /SEEN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104

6.48.9 /SINCE=date-time . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104

6.48.10 /SUBJECT=”string” . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104

6.48.11 /TITLE=”string” . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104

6.48.12 /UNREAD . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104

6.48.13 /UNSEEN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 105

6.49 SEARCH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106

6.49.1 /NEWSGROUPS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106

6.49.2 /EDIT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106

6.49.3 /HEADER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107

6.49.4 /NODISPLAY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107

6.50 SCREEN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108

6.51 SELECT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109

6.51.1 /NEW . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109

6.51.2 /MARK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109

6.52 SEND . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110

6.52.1 /AUTHOR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110

6.52.2 /EDIT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110

6.52.3 /EXTRACT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110

6.52.4 /HEADERS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110

6.52.5 /QUOTE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 6.52.6 /MEMBERS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111

6.52.7 /SELF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111

6.52.8 /SUBJECT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111

6.52.9 /TO . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111

6.53 SET ACCESS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112

6.54 SET BROADCAST_TRAPPING . . . . . . . . . . . . . . . . . . . . . . . . . . . 113

6.55 SET CLASS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114

6.56 SET DISPLAY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 115

6.56.1 SCANSIZE=n . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 115

6.57 SET HISTORY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 116

6.58 SET ITEM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 117

6.58.1 /HOLD=n . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 117

6.59 SET LINE_EDITING . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 118

6.60 SET MANAGER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119

6.61 SET [NO]MODERATOR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 120

6.62 SET NEWSGROUP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 121

6.62.1 /HOLD=n . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 121

6.62.2 /ITEMHOLD=n . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 122

6.62.3 /MODERATOR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 122

6.62.4 /LOCAL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 122

6.62.5 /RESTRICT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123

6.62.6 /WRITE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123

6.62.7 /SERVER=node . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123

6.62.8 /KEEPREQUEST . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 124

6.62.9 /FILEHOLD=n . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 124

6.62.10 /PROTOCOL=p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 124 6.63 SET PROFILE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 125

6.63.1 /AUTOMATIC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 125

6.63.2 /CLASS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 125

6.63.3 /DIRECTORY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 125

6.63.4 /EDITOR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 126

6.63.5 /PERSONAL_NAME . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 126

6.63.6 /PRINT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 127

6.63.7 /TEMPORARY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 127

6.64 SET SEEN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 128

6.64.1 /POINTER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 128

6.64.2 /BEFORE=date . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 128

6.64.3 /FOLLOWUP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 128

6.65 SHOW CLASSES . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 129

6.66 SHOW CONFERENCE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 130

6.67 SHOW ENTRY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 131

6.68 SHOW ITEM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 132

6.69 SHOW KILL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 133

6.70 SHOW MARKER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 134

6.71 SHOW MEMBER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135

6.71.1 /ALL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135

6.71.2 /MODERATOR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135

6.72 SHOW MODERATOR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 136

6.73 SHOW NEWSGROUP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 137

6.74 SHOW NOTE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 138

6.75 SHOW PROFILE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 139

6.76 SHOW VERSION . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 140 6.77 SKIM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 141

6.77.1 /DIRECTORIES . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 141

6.77.2 /FILES . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 141

6.77.3 /HISTORY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 141

6.77.4 /ITEMS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 142

6.77.5 /NEWSGROUPS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 142

6.77.6 /SERVED . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 142

6.77.7 /ARCHIVE=filename . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 142

6.78 SKIP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 143

6.78.1 /ALL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 143

6.78.2 /POINTER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 143

6.78.3 /BEFORE=date . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 143

6.78.4 /FOLLOWUP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 143

6.78.5 /NEWSGROUP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 144

6.79 SPAWN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 145

6.80 TOPIC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 146

6.80.1 /HEADER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 146

6.80.2 /EDITOR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 146

6.80.3 /ROT13 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 146

6.81 UNMARK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 147

6.82 UNREAD . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 148

6.83 UP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 149

6.84 UPDATE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150

6.85 VERSION . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 151

6.86 WRITE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 152

6.86.1 /EDIT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 152 6.86.2 /CONTROL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 152 6.86.3 /HEADERS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 152 6.86.4 /NEWSGROUPS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153 6.86.5 /SUBJECT=subj . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153 6.86.6 /SELF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153

7 Moderated Newsgroups 154 7.1 Management of Moderated Newsgroups . . . . . . . . . . . . . . . . . . . . . . . 154 7.2 POST . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 155 7.3 ADD FILE/MODERATOR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 155 7.4 CANCEL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 155

8 ( 156 8.1 NEWSGROUP MANAGEMENT . . . . . . . . . . . . . . . . . . . . . . . . . . . 156 8.2 ( . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156

9 NEWS Distribution 158 9.1 SRC distribution . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 158 9.2 OBJ distribution . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 160 9.3 DIST distribution . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 160 9.4 DOC distribution . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 162

10 Installing NEWS 163 10.1 Installing NEWS files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 163 10.2 Compiling and Linking NEWS . . . . . . . . . . . . . . . . . . . . . . . . . . . . 163 10.3 NEWS Accounts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 164 10.4 Logical Names . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 164 10.5 Global DCL Symbols . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 166 10.6 Install NEWS.EXE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 166 10.7 Initial Environment . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 166 10.8 NEWS Control Files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 167 10.9 First Time Execution of NEWS . . . . . . . . . . . . . . . . . . . . . . . . . . . . 167

11 Upgrading from Previous NEWS Versions 169

12 Installing NNTP NEWS Servers 170

12.1 DECNET Server and Client Nodes . . . . . . . . . . . . . . . . . . . . . . . . . . 170

12.1.1 Compatibility with Previous NEWS Versions . . . . . . . . . . . . . . . . 171 12.2 CMU TCP Server and Server/Client Nodes . . . . . . . . . . . . . . . . . . . . . 172 12.3 WIN/TCP Server and Server/Client Nodes . . . . . . . . . . . . . . . . . . . . . 172

13 NEWS Feed to a Unix Site 173 13.1 NNTP-based NEWS feeds . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173 13.2 Mail-based NEWS feeds . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 174

Chapter 1

Abstract

This document describes the NEWS program written for VMS. It is intended for users of NEWS and also for system and network managers in order to configure the NEWS system in the local host system.

The first chapter is a general introduction to the function of news programs. The second part is a user manual for NEWS. The third part is a system manager’s guide to NEWS, covering maintenance issues and automatic procedures. The last section is also a system manager’s guide, describing how the NEWS program is configured on a VMS system and also into a wider network of systems.

1.1 Author

 Author:     Geoff Huston
 Address:    Computer Services Centre
             Australian National University
             G.P.O. Box 4, Canberra City, A.C.T.  2601
             AUSTRALIA
             (062)493385
 ANU-DECNET:     csc::gih900
 ACSnet,CSNET:   gih900@csc.anu.oz
 INTERNET:       gih900%csc.anu.oz@uunet.uu.net
 UUCP:           {uunet,ubc-cs,nttlab,mcvax,ukc}
                     !munnari!csc.anu.oz!gih900
 VAX_PSI:        PSI%050526244003211::gih900

Introduction

NEWS is a computer-based conferencing system. As such, it has many similarities to a normal electronic mail system (e-mail), as well as some major differences:

• E-mail is typically a person-to-person communication. In such a system a third party cannot view the mail contents, and the mail is essentially a transaction between the sender and recipient.

This e-mail model can be extended with the use of mailing lists. In this situation each user posts a mail item to a list of others, each of whom receive a copy of the mail. In this model the mail transaction is extended to include a group of users, but privacy is still intact to the extent that users who are not on the mail distribution list cannot participate in the mail conversation. The mailing list can be maintained on a central site, or each user may keep a copy of the list - the resultant functionality from the user point of view is essentially unchanged.

• Bulletin boards take the mailing list model one step further by allowing all users to have the ability to view all postings, and all postings made by users are public to all other users of the bulletin board (this is a simplification of many bulletin boards, which commonly provide functionality which encompasses all these models of communication, and more). As with mailing lists, the bulletin board data can be held on a single central system (acting as a server for client systems), or each system in a network may hold a copy of the postings. In both cases the functionality is similar from the user viewpoint.

VMS NEWS is an implementation of such a bulletin board. VMS NEWS allows the user to view all postings made by other users of NEWS, and also allows the user to contribute to the conference by posting an item into NEWS. News items are classified by into newsgroups (postings with similar subject matter) and all postings referring to the same conference are logically grouped together for presentation within a newsgroup (this is analogous to VMS Mail’s use of folders).

Unlike person-to-person mail, all items posted into NEWS are (normally) readable by ALL users on the system, so the items in news are intended for general view. Typically, the local VMS system is linked to a wider network of systems, and the local node will receive NEWS postings from other nodes on the network, and local postings will also be sent through the wider network. This default public status of news postings is qualified by the ability to configure ”closed” newsgroups with an explicit readership list.

NEWS implements the Standard for Interchange of USENET Messages (RFC1036), and as such has many similarities with the Unix(tm)-based rnews software set. However NEWS is not a port of this rnews software to VMS and the program has been designed with the aim of implementing the functionality of rnews using data and file structures that can be executed efficiently on a VMS system.

The implementation of this standard under VMS implies that a VMS site can be configured into the USENET news network, and will accept USENET news batches from neighbouring USENET sites, and generate batches for transmission to downstream sites.

NEWS also implements the Network News Transfer Protocol (NNTP), as documented in RFC977. This implementation uses DECNET as the transport layer, and is implemented as a DECNET object. The NEWS program can be configured to automatically access a NNTP server over DECNET, implementing a distributed news database. This version also supports the CMU TCP/IP package for VMS, and allows NNTP client and server implementations on VMS accessing remote VMS or Unix systems over TCP.

2.1 Sources of Further Information

• The HELP command is available within NEWS for help on all NEWS commands.

• The information contained in the newsgroups within NEWS are particularly useful: in particular the items in newsgroup ”news.announce.newusers” should be read, as there are postings in this newsgroups which will detail much of the structure of the USENET news network. Users should read these items, particularly those referring to the rules for posting into the USENET and the network etiquette of posting.

The User Interface to NEWS

3.1 Invoking NEWS

To invoke an interactive session of NEWS, use the command:

    $ NEWS

NEWS can operate in either line or screen mode. Line mode is intended for use on printing terminals or within batch command procedures and the user interface is more restrictive than screen mode. The command ”$ NEWS” invokes screen mode, and the command ”$ NEWS/NOSCREEN” starts NEWS in line mode. (You can switch between line and screen modes within NEWS using the SCREEN and NOSCREEN commands described below).

NEWS uses the VMS SMG routines as the screen display manager. If your terminal type is not defined to SMG then NEWS will automatically enter line mode, and the ’SCREEN’ command is disabled.

When starting NEWS your context will be restored to the state when you last EXITed from NEWS. NEWS will re-establish the previous ’current’ newsgroup and newsitem if possible.

To check if there are new unseen messages in the news area additional qualifiers can be used:

    $ NEWS/SCAN

will check the status of the news database and will output a 1 line message if there are unseen items in the news database. The check of newsgroups only includes those newsgroups for which you want the NEWS software to check - the section in the following chapter on keeping track of newsitems indicates how this is done. If there are no unseen items, NEWS will exit without printing a message.

This command may be placed in your LOGIN.COM to indicate that there is unseen news in a format consistent with VMS’s new mail message.

A more verbose form of this scan can be achieved with the command:

    $ NEWS/ALLSCAN

The differences are that this qualifier directs NEWS to list those newsgroups which contain unseen items, and the number of such items in each newsgroup. Also if there are no unseen items then a status message is output indicating this. Again this command may be placed in your LOGIN.COM file.

NOTE that these two command qualifiers do not execute rapidly - the execution time for these two commands is proportional to the number of newsgroups carried on your local NEWS database.

You can also invoke NEWS on a conditional basis: if there are unseen items then NEWS will start an interactive NEWS session, and if there are no unseen items NEWS will immediately exit. The command format for this action is:

    $ NEWS/UNSEEN

3.2 NEWS Context

When in NEWS you maintain an implicit ’context’ within the NEWS database. This context is made up of three components:

• The ’current’ newsgroup is the newsgroup indicated by the cursor (when at the newsgroup directory level). This is the newsgroup which is ’opened’ if you are at the newsitem directory or text display levels.

• The ’current’ newsitem - which is relevant only if the screen display is either the newsitem directory (in which case the ’current’ newsitem is indicated by the cursor) or newsitem text display.

• The screen level - either a newsgroup directory, a directory of items within a newsgroup, or the text of a newsitem.

3.3 The Structure of NEWS

NEWS items are stored within NEWSGROUPS (this is analogous to VMS MAIL, where mail items are stored within FOLDERS). All news items with a similar subject matter are normally contained in a single newsgroup.

Each news item within a newsgroup has a title and a sequence number. A directory of items within a newsgroup displays these titles and sequence numbers - the sequence number is used to select a particular item from within a newsgroup. News items contain other control information - these additional fields will be covered in the section on Posting news items.

3.4 NEWS Screen Formats

There are three different screen display formats used in NEWS. They correspond to the context level of the user when running NEWS (Newsgroup directory, Newsitem directory and Newsitem text).

The top three lines on each display describe the current context level, and any other necessary information for that screen type.

At the base of all three screen formats is a three line window. The first line shows a command prompt – commands are entered on this line. The second line is used to enter any additional parameters required for the command. The third line is used to display any status or error messages (in reverse video).

3.4.1 NEWSGROUP Directory Screen

This screen display corresponds to the outer context level of NEWS. The screen displays the names of the newsgroups held on the system, the number of news items held in each newsgroup and the number of items that you have not read.

The top line displays the selection criteria of newsgroups: this may be all newsgroups, newsgroups in which you are registered, newsgroups with new items posted into them since a certain date, and so on. The line also displays the number of newsgroups which have been selected under these selection criteria.

The major section of the screen displays the newsgroup names, their sequence number, and additional control information.

The following figure explains the various fields shown on the directory screen:

  Number     Newsgroup                           Count Unread
 -->  45  _comp.os.vms                            23     12
  |   |   |  |                                       |      |
  |   |   |  The name of the group    The total number      |
  |   |   |                           of newsitems  in      |
  |   |   If this character is a '_'  the newsgroup         |
  |   |   then  some  items  in  the                        |
  |   |   newsgroup  are stored on a            The number of
  |   |   remote news server node               newsitems you
  |   |                                         have not seen
  |   The sequence number of the newsgroup
  |
  The 'current' newsgroup pointer

From left to right, the fields are: the cursor arrow, used to select the current newsgroup (this arrow is moved using the up and down arrows on the keypad), the newsgroup sequence number, a ’_’ character (only displayed if this newsgroup is linked to a remote news server), the newsgroup name, the count of items in the newsgroup, and the number of unread items in the newsgroup.

Newsgroup names in which you have registered an interest are shown in bolded video on the screen.

In the leftmost field a cursor is displayed (–>). Pressing the up or down arrow keys moves the cursor through the newsgroup display. The cursor indicates the current or selected newsgroup.

3.4.2 NEWSITEM Directory Screen

The second context level is entered from the NEWSGROUP screen by pressing the select key (or using the ’SELECT’ command). This screen corresponds to the second context level, when a newsgroup is ’opened’ to display the news items currently stored within the newsgroup. An example screen is shown below:

 Newsgroup: anu.news  (1 items: #6 - #6)
 Expry: [10], Mod: anu-news@fac3.anu.oz, Loc, [Srv: FAC3::, File: 2 ]
          Title                                           Lines Date
--> 6   _Weekly USENET Report, period ending Mon Feb 15     78 16-Feb-88

These fields are:

 Newsgroup: anu.news  (1 items: #6 - #6)
            |          |
    Newsgroup name   Number of items, and first and last
                     item sequence numbers

  Expry: [10], Mod: anu-news@fac3.anu.oz, Loc, [Srv: FAC3::, File: 2 ]
  |            |                          |     |            |
  |            Internet address of        |  Decnet  address |
  |            newsgroup moderator        |  of server  node |
  |            [not shown if group        |  [not  shown  if |
  |            is not moderated]          |  newsgroup isn't |
  |                                       |  remote served]  |
  Retention time (days) of item           |                  |
  on local system                 Local newsgroup [not  Retention period
                                  shown   for  network  (days) for local
                                  wide newsgroups]      copy  of  remote
                                                        held newsitem

            Title                                         Lines Date

 --> 6   _Weekly USENET Report, period ending Mon Feb 15    78 16-Feb-88
 |   |   | |                                                 |  |
 |   |   | The title of the item       The size (lines) of the  |
 |   |   |                             newsitem text.           |
 |   |   If this character is a '_'                             |
 |   |   then this newsitem is held        The date the item was added
 |   |   on the  remote news server        into the newsgroup.
 |   |   node
 |   |
 |   The sequence number of the newsgroup
 |
 The 'current' newsitem pointer

This display is similar to the newsgroup display, using 1 line to describe each newsitem. The screen shows the sequence number, news item title, date the news item was created on the local system, and the size of the news item. News items which you have not viewed are shown in bolded video. As with the newsgroup directory, a cursor is positioned in the leftmost columns of the screen, and the cursor can be moved within the screen by the up and down arrow keys. The cursor corresponds to the current or selected news item.

3.4.3 NEWS ITEM Text Display

The third context level is entered by the ’READ’ command. This screen corresponds to the third context level, displaying the contents of the current news item.

The text is displayed in screen sized ’chunks’, advancing through the text one page at a time. The base of the screen shows the current position within the text as a percentage of the total text size. The up and down arrow keys are also mapped in this mode, and move the display in single line movements in the relevant direction. The Prev-Screen and Next-Screen are also mapped to move the screen in page sized chunks either up or down. Lastly the commands UP and DOWN with a numeric parameter (e.g. DOWN 150) will move the specified number of lines. The underlying display buffer is 500 lines in size, and once text rolls off the ”top” of this display it cannot be presented back on the screen.

Optionally you may use an editor (TPU by default, but you may choose another editor) in read-only mode to display the news item text. This will allow you to scroll up and down, search for strings, and use other editor search and display operations on the text. At any stage the text display can be aborted, and you may reset the screen display to either the news item or newsgroup display screen.

(

Basic NEWS Commands)

This is the first of three chapters describing the user-level NEWS commands. This chapter describes those commands which are commonly used, and form the minimum set of commands necessary to use NEWS. Chapter 5 describes those commands which are used to change the NEWS environment, and chapter 6 is a reference chapter for the user command set, describing in detail all user commands and their qualifiers.

As with VMS DCL commands, all NEWS commands and their qualifiers may be abbreviated to the minimum unique abbreviation.

4.1 Default Keypad Definitions

By default, the keypad is configured in the following way:

+--------+--------+--------+--------+
|  pf1   |  pf2   |  pf3   |  pf4   |
gold|        |        |        |SKIP/FOLLOWUP
|GOLD    |HELP    |SKIP    |SKIP/NEWSGROUP
+--------+--------+--------+--------+
|   7    |   8    |   9    |   -    |
|PRINT   |EXTRACT |EXTR/ALL|EXTR/APPEND
|DIR     |DIR/NEW |DIR/REG |DIR/ALL |
+--------+--------+--------+--------+
|   4    |   5    |   6    |   ,    |
|        |TOP     |        |        |
|DIR     |>>UP<<  |DIR/REG |DIR/ALL |
+--------+--------+--------+--------+
|   1    |   2    |   3    | enter  |
|READ/HEA|BOTTOM  |READ/NEX/HEAD    |
|READ    |>>DOWN<<|READ/NEXT        |
+--------+--------+--------+        |
|   0             |   .    |        |

|READ/NEW/HEADER  |READ/NEW/FOLLOWUP/HEADER
|READ/NEW         |READ/NEW/FOLLOWUP|
+-----------------+--------+--------+
   VT100/VT220 KEYPAD BLOCK

+--------+--------+--------+
|  find  | insert | remove |
|        |        |        |
| SEARCH |REGISTER|DEREGISTER
+--------+--------+--------+
| select |  prev  |  next  |
gold|        | TOP    |BOTTOM  |
| SELECT | UP 18  |DOWN 18 |
+--------+--------+--------+
VT220 EDIT KEY BLOCK

These default key mapping may be altered by the user - this is documented in chapter 5.

4.2 NEWS Screen Displays

This set of commands enable you to move around the NEWS database.

DIRECTORY The default action of NEWS is to display the newsgroup directory screen. This screen has a cursor in the left field, which may be moved using the arrow keys. The directory of newsitems within the ’pointed’ or ’current’ newsgroup may be obtained by the ’DIR/ITEMS’ command.

When the item screen display is displayed, the newsgroup directory may be displayed using the ’DIR/NEWSGROUPS’ command.

The set of newsgroups displayed in the newsgroup directory may be modified using qualifiers to the DIRECTORY command: DIR/ALL displays all newsgroups, DIR/REGISTER displays only those newsgroups in which you have registered a particular interest, and DIR/SINCE=date (e.g.: DIR/SINCE=TODAY) displays only those newsgroups which have received new postings since the data specified.

UP, DOWN The newsgroup and newsitem directory screens use a pointer in the left field of the display to indicate the ’current’ newsgroup and newsitem. This pointer may be moved up or down using the UP or DOWN commands. These commands are bound to the arrow keys, so the more usual way of moving through the screen is via the arrow keys.

Both commands may take a numeric parameter (e.g.: UP 18) to move a number of lines. On VT220-type terminals the PREV SCREEN and NEXT SCREEN are bound to the commands ’UP 18’ and ’DOWN 18’ respectively.

SELECT is similar to ’DIR/ITEMS’ such that the command displays a newsitem directory of a newsgroup. However, SELECT has a number of qualifiers which make the selection of newsgroups more powerful.

In particular, ’SELECT/NEW’ selects the next newsgroup in which you have registered a particular interest which has received new newsitems.

4.3 Displaying NEWS text

READ The most common method of displaying the contents of a newsitem is using the ’READ’ command. This displays the news text on a page-by-page basis (in a similar fashion to the DCL ’TYPE/PAGE’ command).

By default, the ’current’ newsitem is opened for display, but qualifiers may be used to combine movement of the ’current’ newsitem and display: READ/NEXT combines the actions of the ’DOWN’ and ’READ’ commands. READ/NEW combines the actions of ’SELECT/NEW’ (select the next unread newsitem as the current item) and ’READ’. READ/PARENT selects the parent newsitem (the item to which the current newsitem refers) for display. READ/FOLLOWUP selects the next followup newsitem (a posting which refers to the same subject as the current item) for display.

4.4 Posting Items into NEWS

There are two commands which allow the user to create a NEWS posting: POST and FOL-LOWUP. POST implies that the item refers to a new subject which has not been discussed before: FOLLOWUP implies that the user is posting a followup item to the current item.

The default actions of POST and FOLLOWUP are to forward the item text as far as possible throughout the NEWS network. Background NEWS server processes are responsible for sending the item to all adjacent network neighbours. Normally the distribution of the item is restricted by the set of network nodes which wish to receive the newsgroups in which the newsitem is posted. Qualifiers may be used with these commands to restrict this distribution to a subset of the network (these qualifiers are described in the reference chapter, Chapter 6).

POST invokes a text editor to create a new newsitem (the text editor may be one of your choosing - see the chapter on customization). The NEWS system also prompts for the newsgroup or newsgroup list to which the item will be posted, and will also prompt for the title of the posting.

FOLLOWUP is similar to POST, with the difference that the edit buffer is pre-loaded with the text of the current newsitem (the item to which you are posting the followup). This allows you to formulate your reply using fragments of the original posting to help other readers establish what you are replying to.

4.5 Deleting Newsitems

Deletion of newsitems is not normally a straightforward process - as the text of the original item may have already been forwarded to the wider news network, then deletion of a newsitem requires that a special deletion control message must be sent through the network to delete all copies of your posting.

CANCEL This command deletes the ’current’ newsitem. A check is performed to ensure that the item you are attempting to delete was actually posted by yourself before the deletion is performed. This command also generates a cancel control message which is forwarded throughout the net, which directs all remote nodes to delete this newsitems on your behalf.

4.6 NEWS Interface to MAIL

NEWS supports two commands which interface to VMS MAIL. This allows users to forward NEWS items to other users via MAIL and also to send direct person-to-person mail to the original poster of the newsitem.

FORWARD, MAIL These two commands are synonyms: the ’current’ newsitem is extracted and posted to the addresses you supply. By default the item text is loaded into an edit buffer for editing before sending the mail - this edit pass may be omitted by using the ’/NOEDIT’ qualifier with the command.

REPLY This command extracts the From: address (or the Reply-To: address, if specified) from the current newsitem and invokes the editor to create a reply which will be directed to the sender of the newsitem.

This command differs from the FOLLOWUP command in so far as your response to the original posting is a private response between you and the poster, whereas the FOLLOWUP command posts a public response through the NEWS network.

4.7 Extraction of Newsitems

Newsitems may be extracted from the NEWS database and kept as a local file, or printed.

EXTRACT This command extracts the ’current’ newsitem and makes a local copy into a file - you are prompted for a filename to write to.

PRINT This prints the ’current’ newsitem to a print queue.

4.8 Keeping Track of Newsitems

There are two ways in which NEWS can be used to help you read particular newsitems - the first is by the concept of ’read’ and ’unread’ newsitems and the second by the use of ’marks’ to enable you to re-read newsitems at a later date.

NEWS maintains for each user a list of all newsitems which have been considered as ’read’ - this is an aid when searching through the NEWS database for new items which have been posted. NEWS also maintains for each user a list of newsitems which have been ’marked’ by that user, enabling the user to ’mark’ a newsitem with any tag text, and later retrieve that newsitem by specifying that tag text.

This context information is contained in a personal file, ”SYS$LOGIN:NEWSRC.;”. The file is updated every time NEWS is executed.

REGISTER, DEREGISTER NEWS enables the user to mark particular newsgroups as being of particular interest. This marking, or ’registering’, is used by the READ/NEW, SELECT/NEW, DIR/NEW and DIR/REGISTER commands to allow fast access to any new postings in the subset of ’registered’ newsgroups.

The ’current’ newsgroup is added to this registered set by the ’REGISTER’ command, and can be removed from the set by the ’DEREGISTER’ command.

SKIP is used to mark a set of newsitems as having been ’read’ without having to display the item text. The default action is to mark the current item as ’read’ and move the pointer to refer to the next item.

If the display is the newsgroup directory, then SKIP will mark ALL newsitems in the ’current’ newsgroup as read.

SKIP may also mark a range of newsitems as ’read’ by using qualifiers with the SKIP command - these are documented in the reference chapter.

MARK, UNMARK are used to associate tags with a newsitem. The common form of marking is the MARK the ’current’ newsitem, and then retrieve that item at a later date by either the SELECT/MARK or READ/MARK commands. If a newsitem has been marked, then a line of the form ”Mark: ...” is included in the header lines of the text display for that newsitem.

A mark may be removed from a newsitem by the UNMARK command.

Variants of this command allow the mark to have a text value, allowing the READ/MARK and SELECT/MARK commands to retrieve particular marked items according to a match on the supplied tag values.

4.9 HELP

HELP is used to enter the on-line HELP documentation - this documentation is similar to the reference chapter in this document

4.10 EXITing NEWS

EXIT performs a normal exit from NEWS - the register file is updated to reflect all items read during this session, and any pending print requests are sent to a nominated print queue.

QUIT exits from NEWS without the associated update function of EXIT. This is a more graceful alternative to the Ctrl-Y exit.

(

Customising the NEWS Environment)

It is possible to customise NEWS in a number of ways. This includes changing the definitions of keys on the terminal, and changing the editor invoked by NEWS to one of your choice. This chapter describes those commands and logical names which perform this customisation. Customization can be done using three mechanisms: setting your keypad, as described below, using the SET PROFILE command, and defining logical names. As there is some overlap in what these three methods can define the order of precedence is firstly any setting performed interactively overrides profile settings, which in turn override logical name settings.

5.1 Key Definition

The default set of key mappings is given in the previous chapter.

These defaults may be altered in two ways: using the ’DEFINE/KEY’ command to bind a news command to a key, or defining the logical name ”NEWS_INI” to a file containing a sequence of DEFINE/KEY lines (which will be executed on startup of NEWS).

The DEFINE/KEY command is equivalent to it’s VMS DCL counterpart - the command performs a mapping of the nominated key to a command string.

NOTE: There are six keys which should not be DEFINED from the standard NEWS values: the four arrow keys and the PREV and NEXT SCREEN keys (This restriction is a reflection of an underlying problem with the VMS software used by NEWS).

5.2 Logical Names

There are a number of logical names which users may define to alter the behaviour of NEWS:

NEWS_INI This can be defined as a file name which contains a set of DEFINE/KEY commands to change the default key binding of news. e.g.

  $ DEFINE NEWS_INI SYS$LOGIN:NEWSKPD.DEFS

NEWS_EDIT This logical name can be used to link an editor to NEWS. The editor is used to generate new postings, and to edit forwarded mail, and (optionally) in readonly mode, to scan the item contents (READ/EDITOR).

By default NEWS links the TPU editor to NEWS, using the value of the logical name TPUSECINI as the section to use. With V5 of VMS the NEWS software will also examine whether the logical name TPU$SECTION is defined and use that translation in a similar fashion to TPUSECINI.

This logical name is used to either:

• Use the TPU editor with a section other than the default:

      $ DEFINE NEWS_EDIT "TPU/SEC=mysection"

• Use EDT as the configured editor:

      $ DEFINE NEWS_EDIT CALLABLE_EDT

• Call any other editor using a command procedure in a subprocess:

      $ DEFINE NEWS_EDIT "@dev:[dir]my_editor.com"

and NEWS will then SPAWN the command:

      $ @dev:[dir]my_editor.com filename

The next section shows an example my_editor.com command procedure.

NEWS_VIEW This logical name is only required if you are using a local editor command procedure as NEWS_EDIT. In this case a second procedure is required which invokes the same editor in read_only mode:

  $ DEFINE NEWS_VIEW "@dev:[dir]my_reader.com"

and NEWS will then spawn the command:

  $ @dev:[dir]my_reader.com filename

When TPU or EDT is configured as the editor, then NEWS will automatically add the /READONLY to create a readonly version of the editor.

NEWS_SCANSIZE NEWS will normally display the text of a newsitem on a page-by-page basis. When scanning through items it is often useful to only display the first few lines of an item, to decide whether to view the complete text.

If the logical name NEWS_SCANSIZE is defined as a number (n), then the first n lines of each new item will be displayed, and you will be prompted if you wish to see more of this item. Commonly, a value between 8 and 12 is most useful.

MAIL_SIGNATURE This logical name may be defined as the name of your mail signature file. If this logical name is defined, then the contents of this fill will be automatically appended to all postings and mailings made from within NEWS.

NEWS_DEFAULT_DIRECTORY NEWS will normally remember the last newsgroup directory setting across invocations of the program (i.e. If you had executed DIR/NEW before exiting NEWS, then when you next run NEWS, the initial newsgroup directory state will be restored to DIR/NEW). It is possible to specify a constant initial directory state by defining this logical name as one of the strings:

     ALL  NEW  REGISTER  SINCE  SINCE=vms_date

NEWS_POST_CC This logical name directs NEWS to generate MAIL carbon- copies of NEWS postings. The logical name value is interpreted by NEWS as the name of a CC control file. This file consists of lines of the form:

    newsgroup   address

The newsgroup name starts in column 1, and may be the name of a single newsgroup, or may specify a set of newsgroups by including the wildcard match character (’*’).

The address is the address to receive a carbon copy of the posting. It is separated from the newsgroup name by at least 1 space character. The address may be a decnet or internet-format address.

The interpretation of the file is such that any postings you make into any of the specified newsgroups will also generate mail to the matching addresses.

NEWS_PERSONAL_NAME Sets the personal name string. This name is appended to your network address in all NEWS postings.

If this logical name is not defined your VMS MAIL profile is examined to extract your MAIL personal name as your NEWS personal name.

5.2.1 NEWS_EDIT Command Procedure

An example command procedure to invoke an editor from NEWS is as follows:

$ !
$ ! Example command procedure to invoke an editor for NEWS
$ !
$ ! Inputs:
$ !
$ !     P1 = file name.
$ !
$ ! Note that this procedure is run in the context of a subprocess.
$ ! LOGIN.COM is not executed.  However, all process logical names
$ ! and DCL global symbols are copied.
$ !
$ ! The default directory is the same as the parent process
$ !
$ ! Note that this will only be invoked if the user has placed in
$ ! LOGIN.COM (or similar)
$ !     $ DEFINE NEWS_EDIT "@dev:[dir]NEWSEDIT_EXAMPLE.COM"
$ !
$ DEFINE /USER SYS$INPUT 'F$TRNLNM("SYS$OUTPUT")'

$ IF P1 .EQS. "" THEN EXIT
$ EDIT/EDT 'P1'
$ EXIT

NEWS Commands - User Reference

All commands in this section are listed alphabetically, not by function. Also all moderation and management commands are included in this section, are are noted as restricted commands.

6.1 $ NEWS

Format:

    $ NEWS

The NEWS command is the VMS DCL command to invoke NEWS.

6.1.1 /SCAN

Format:

    $ NEWS/SCAN

NEWS can be invoked to simply check if there are unseen items in the NEWS area, then exit to DCL. The qualifier directs NEWS to check the status of the news database and will output a 1 line message if there are unseen items in the news database. The check of newsgroups only includes those newsgroups for which you want the NEWS software to check - registered newsgroups. If there are no unseen items, NEWS will exit without printing a message.

This command may be placed in your LOGIN.COM to write out a message that there is unseen news, in a format consistent with VMS’s new mail message.

6.1.2 /ALLSCAN

Format:

/ALLSCAN is a more verbose form of the /SCAN action. The difference is that this qualifier directs NEWS to list the names of those newsgroups which contain unseen items, and the number of such items in each newsgroup. Also if there are no unseen items then a status message is output indicating this. Again this command may be placed in your LOGIN.COM file.

NOTE that these two command qualifiers (/SCAN and /ALLSCAN) do not execute rapidly - the execution time for these two commands is proportional to the number of newsgroups carried on your local NEWS database.

6.1.3 /UNSEEN

Format:

    $ NEWS/UNSEEN

NEWS can also be invoked on a conditional basis: if there are unseen items then NEWS will start an interactive NEWS session, and if there are no unseen items NEWS will immediately exit.

6.1.4 /SCREEN

Format:

    $ NEWS/SCREEN  [Default]
    $ NEWS/NOSCREEN

This qualifier is used to determine whether the initial state of NEWS is configured to drive a terminal screen or a hardcopy device. The /NOSCREEN qualifier is used for hardcopy interactive sessions or Batch or Detached execution of the NEWS program.

See Also SCREEN and NOSCREEN commands

6.1.5 Parameters

Format:

    $ NEWS command

Any valid NEWS command may be specified as a parameter to the NEWS command: If such a parameter is used then NEWS will execute this single command then exit to the VMS DCL level.

For example, the VMS DCL command:

    $ NEWS POST/NEWSGROUP=local.msg/SUBJECT="Weekly Msgs" weekly.msg

posts the file ’weekly.msg’ into NEWS, into newsgroup ’local.msg’, with the title ”Weekly Msgs”, then returns to DCL.

6.2 ADD ENTRY

Format:

  ADD ENTRY newsgroup

ADD ENTRY is used to add a newsgroup into a set of related newsgroups (or ”class”). This set of newsgroups is defined and maintained by the user, and held in the user NEWSRC file. Classes are a user tool to allow the user to group a number of related newsgroups together in a single area (Classes are very similar to the use of folders in the VMS MAIL utility).

Normally the specified newsgroup is added to the current class (if a class is currently selected). The /CLASSES qualifier is used to specify the set of classes in which the newsgroup is to be entered. Note that if the newsgroup is currently a member of a class (or classes) the classes specified in this command are interpreted as additions to the current class set.

6.2.1 /CLASSES

Format:

    ADD ENTRY/CLASSES=(class-name[,...])

This qualifier determines the classes to which the specified newsgroup is to be added. If this qualifier is not used, then the current class is used as the default value. If a class of the specified name does not exist, it is created by this command.

The qualifier value may be a list of class names, in which case the newsgroup is added to all named classes. The class names may also contain wildcard characters (’%’, ’*’), in which case the newsgroup will be added to all classes with matching names.

See Also SET CLASS, SHOW CLASSES, SHOW ENTRY, MODIFY ENTRY, DELETE EN-TRY, DIRECTORY/CLASS

6.3 ADD FILE

ADD FILE is a NEWS MANAGEMENT command.

Format:

    ADD FILE filename

ADD FILE is used to load articles into the NEWS database. This command is normally used to load network news batches onto the local node (locally generated news is POSTed into NEWS by the user who is executing the POST command).

If the node is defined as a NEWS network routing node, this command also generates the batch files to send to the adjacent NEWS nodes.

ADD FILE scans an input file and breaks the file into one or more articles. Each article is composed of a number of header lines (or a number of header line groups) and an article body. The header lines may be generated by a number of different sources, and a number of these formats may appear at the head of the article. If VMS Mail is used as the local delivery mechanism, then the the article will contain the three line Mail-11 headers. If the item was cross-mailed from an SMTP mailer then the RFC822 header lines may also be included in the article. If the article was generated by a remote news system then there will also be a set of NEWS headers included as well.

The ADD FILE command reads groups of header lines until the first non-valid header line. The resultant set of headers is filtered, and any non-NEWS headers (as defined in the relevant standard) are discarded before the news item header line set is generated. Also local values are inserted for any required NEWS header lines which were omitted in the item.

This implies that VMS Mail folders can be readily read into NEWS - for example:

    MAIL> set folder tcpip
    MAIL> extract/all tcpip.news
    MAIL> exit
    $ NEWS add file tcpip.news/defnews=local.mail_list.tcpip

Each input file may contain one, or a number of news items. The item delimiter is parsed as either of the following character sequences:

    "<newline>#! rnews <size>"   - RFC1036

- or -

    "<formfeed><newline>From: "  - VMS Mail

The item parsing will also automatically recognise the start of an item written using the leading ’N’ format and automatically strip each line before passing it to the parser (This format, which inserts a ’N’ character in column 1 of the NEWS item is used by other NEWS implementations when passing extracted NEWS batches into the Mailer for posting.)

The ADD FILE command takes a single parameter, the name of an input file to parse. The filename specified may be either a single filename, or may contain wildcards (which are expanded, and all files so specified are added).

Lists of filenames are not supported in this version of NEWS, however wildcard characters (’*’ and ’%’) are supported (in the standard VMS fashion).

6.3.1 /CREGRP

The default action of the ADD command is to only accept items into newsgroups which already exist on the system, rejecting all other newsgroups. This qualifier directs NEWS to automatically create any new newsgroups which are specified in the item headers which do not already exist on the local system. Default is /NOCREGRP.

The reason for the default action of /NOCREGRP is that ”checknews” control messages normally flow through the net, controlling which newsgroups are valid for a participating NEWS site. These control messages will auto-configure your node if you are receiving a network news feed.

6.3.2 /DELETE

This qualifier directs NEWS to delete the input file after the contents have been added to NEWS. Default is /NODELETE.

6.3.3 /EXECUTE

Format:

    ADD FILE/NOEXECUTE    [default]
    ADD FILE/EXECUTE= RETAIN | LOCAL | DELETE

The standard for USENET news messages defines a number of ’control’ messages, designed to assist all sites in running a common set of newsgroups across the USENET. Various control messages are designed to delete all occurrences of a particular newsgroup on all systems, create new newsgroups, or compare the ’standard’ list of newsgroups against the local set of newsgroups.

The default action of NEWS is to not to create or delete newsgroups - if a control message indicates that a newsgroup should be created or deleted, a DCL command file to perform the action is sent to address ’USENET’. (This action is acceptable within the relevant standard). This corresponds to the default ”/NOEXECUTE”.

If /EXECUTE is specified, then newsgroup creation messages are executed immediately, without the requirement for approval by the local news manager. In the case of newsgroup deletion (”rmgroup” or a ”checkgroups” message) the default action of /EXECUTE is to set the newsgroups which are to be removed to a 1 week retention period. This default action may be modified by specifying either ”/EXECUTE=LOCAL”, in which case any such newsgroups are set to ”/LO-CAL” status, or ”/EXECUTE=DELETE”, in which case such newsgroups are deleted. In all cases a log of the actions performed will be posted to address ”USENET”.

6.3.4 /DEFNEWSGROUP=newsgroup

By default, if a newsitem does not contain a ”Newsgroups: ” header line, then ”junk” is used as the default value for this header.

It is necessary to be able to change this default action to add a set of related items which no not contain valid NEWS header lines.

A common example example, described in the notes above, is that of VMS Mail folder items being read into a Newsgroup.

6.3.5 /NETFEED

/NETFEED [default]. The ADD command is commonly used to process a network feed of NEWS - this implies that the ADD FILE command should also generate the necessary batch files for subsequent transmission to any downstream sites. /NETFEED instructs the ADD FILE command to use the information in the NEWS.SYS and NEWS.DISTRIBUTION files to generate the batch feed files as required.

/NONETFEED bypasses the generation of a network feed in the ADD command.

6.3.6 /JUNK

/JUNK [default] instructs the ADD FILE command to place any part of the input file which cannot be correctly parsed, or any news item which is filtered by the NEWS.SYS file, or where there are no local newsgroups to hold the item, into the local newsgroup ”junk”. This newsgroup is intended as a short-term storage area where the items can be scanned, and the database altered if necessary to accept such items.

/NOJUNK instructs the ADD FILE command to reject all items which cannot be added to the local NEWS database.

6.3.7 /ACCEPT

/NOACCEPT [default]. By default all items posted to moderated Newsgroups required an ”Approved: ” header line before local NEWS systems can accept the item. Without this header line the item is either junked or rejected.

/ACCEPT instructs the ADD FILE command to accept (on the local system only) unapproved postings to moderated newsgroups.

6.3.8 /RETRY

When a network feed news item is rejected by the local node an attempt is made to place the item into the ”junk” newsgroup. This rejection may take place for a number of reasons, but the primary one is that the NEWS.SYS file has rejected this item.

You may wish to make changes to the SYS file and retry adding the item. Normally such an attempt would be rejected, as the Path: name now contains the local node name, and this is regarded as a loop in the news. /RETRY directs the ADD command to ignore the Path: line when checking the news item for addition into the local node.

6.3.9 /MODERATOR

This qualifier allows a newsgroup moderator to insert items into the moderated newsgroup. This differs from the POST command in so far as the original submitter’s address is preserved. This is documented in more detail in the chapter on moderated newsgroups.

6.4 ADD MARKER

Format:

    ADD MARKER [marker-name [item-number]]

ADD MARKER places a mark against the current newsitem. This item can be selected at a later stage by the READ/MARK or SELECT/MARK commands, or printed by the PRINT/MARK command.

If no marker-name value is specified then the default value, ”mark” is used. A marker name is a string. If no item number is specified, the mark is placed against the current item.

See Also MARK, UNMARK, CLEAR MARK, DELETE MARKER

6.5 ADD MEMBER

ADD MEMBER is a NEWS MANAGEMENT command.

Format:

    ADD MEMBER username

ADD MEMBER adds a new user to the membership list of a restricted access newsgroup. Membership of the newsgroup grants the username read access to the newsgroup, and by default also grants write access.

The username may include wildcard characters (’%’,’*’) to include a set or related usernames, or may be an identifier name. If either of these specifications are used it is advised not to include the /MAIL or /NODES parameters to the command.

The newsgroup must have been set as a restricted newsgroup for this command to be executed, and the user must be either the local News Manager, or the moderator of this newsgroup.

6.5.1 /MAIL=mail-address

Specifies a mail address for this member. The address format may be either a VMS MAIL or an internet format address.

6.5.2 /PRIVILEGES=priv-list

Format:

    ADD MEMBER user/PRIVILEGES=(MODERATE,WRITE,NOWRITE)

Associates privileges with this username.

MODERATE allows the user to delete any newsitem, add and delete members of the restricted newsgroup and unconditionally write to the newsgroup.

WRITE implies explicit WRITE access to the NEWSGROUP. This is the default for all newsgroup members unless the newsgroup is SET to /NOWRITE, in which case explicit write access or moderate priv is required.

NOWRITE implies read-only access to this newsgroup.

6.6 ANSWER

ANSWER is used to post a reply to a news item back into NEWS. The followup item will reference the news item you are following up as the parent of the new posting. The command FOLLOWUP is equivalent to ANSWER.

The editor is invoked to form the new posting, with the edit buffer preloaded with the text of the original item. The default editor is the TPU editor, but this is user definable - see the section on customisation to specify a different editor to use.

NEWS will fill in all defaulting news headers, and will prompt for all header items which are non-defaulting (the subject line and the newsgroups to post the item into).

If the newsgroups for the posting are moderated newsgroups, the posting is automatically redirected as mail to the moderator address, and NOT posted back to NEWS.

See Also FOLLOWUP, REPLY, POST

6.6.1 /HEADERS

Answer will normally prompt for the subject of the posting, and a list of newsgroups to post the item into. All other header lines of the news item text will be filled with null values.

The /HEADERS qualifier will force prompting for optional components of the item header. These fields are:

Summary: A one line summary of the contents of the posting.

Reply-To: The (internet) mail address for REPLYing to the item - used if replies are to be sent to a mail address OTHER than yourself (as the originator of the item.

Distribution: A (comma separated) list of keywords designed to limit the distribution of a news item to a subset of the network. These keywords are site specific, but commonly include the keywords ”local” (do not forward to the net) and ”world” (forward as far as possible through the net).

Followup-To: The list of newsgroups where followup postings are to be sent. This is used to move a discussion to another newsgroup.

Keywords: A few relevant, pithy keywords which should indicate to other readers if the item is of interest.

6.7 ATTACH

Format:

    ATTACH [process-name]

Attaches your terminal to the parent process, or a specified sub-process. This command takes one optional parameter, a process name. If no process name is specified, the terminal is attached to the parent process of the current process. See Also SPAWN

6.8 BACK

BACK reads the previous item in the current newsgroup. This is equivalent to the command READ/BACK. The command may be modified by specifying one of the parameters listed below.

6.8.1 /HEADER

Directs NEWS to display all header lines of the newsitem. By default only the subject line, sender address and date of the item are shown in the display.

6.8.2 /EDITOR

Use a callable editor in read-only mode to display the newsitem. By default this is the TPU editor, but this may be altered by the NEWS_VIEW logical name. See the chapter on customizing NEWS for details. For compatibility with previous versions of NEWS, /TPU is equivalent to /EDITOR.

6.8.3 /ROT13

Some newsitem are posted using rot13 encryption. This qualifier decrypts the newsitem for display.

6.9 BACK REPLY

Format:

    BACK REPLY

BACK REPLY is equivalent to READ/PARENT, displaying the previous entry in the same conversation stream as the current item. All command qualifiers for the BACK command may be used with the BACK REPLY command.

6.10 BACK NOTE

Format:

    BACK NOTE

BACK NOTE reads the previous item in the current newsgroup. This is equivalent to the command READ/BACK. All command qualifiers for the BACK command may be used with the BACK NOTE command.

6.11 CANCEL

This command deletes the current newsitem (the directory must be at the newsitem level for this command).

This command inspects the address of the sender of the newsitem, and will only perform the deletion if the sender’s address matches that of the user. In the case that the user is a local newsgroup manager, or the local news manager, such a user can CANCEL any posting originating from the current node.

As well as deleting the local copy of the newsitem, a ’cancel’ control news message is sent through the news network to ensure that all copies of the item, held on all connected NEWS nodes, are deleted.

6.12 CLEAR KILL

Format:

    CLEAR KILL
    CLEAR KILL tag

CLEAR KILL removes the kill filters with the matching tag from the kill filter set. If no tag value is specified then ALL kill filters are cleared. See Also KILL, SHOW KILL, MODIFY KILL

6.13 CLEAR MARK

Format:

    CLEAR MARK
    CLEAR MARK tag

CLEAR MARK removes all marks of the specified tag from the mark list. If no tag value is specified then ALL marks are cleared. Otherwise only the specified tag value is cleared. The wildcard character (’*’) can be used in the tag field to clear a set of related tags. See Also ADD MARKER, MARK, UNMARK

6.14 CLOSE

CLOSE closes the current newsgroup, and the display is reset to the directory list of newsgroups.

6.15 CREATE NEWSGROUPS

CREATE NEWSGROUPS is a NEWS MANAGEMENT command.

There are two equivalent commands to this command: CREATE CONFERENCE and CREATE GROUPS.

Format:

    CREATE NEWSGROUPS newsgroup

CREATE NEWSGROUPS is used to create new newsgroups on the local node.

The Newsgroup parameter is the name of the newsgroup to create. If the newsgroup does not exist on the local node, then the user is prompted to confirm that the newsgroup should be created.

This parameter can also specify a comma separated list of newsgroup names to create. In the case of a list of names the qualifiers used (if any) refer to ALL newsgroups in the list (positional qualifiers are NOT supported in this release).

6.15.1 /HOLD=n

This qualifier specifies the retention period (in days) of the Newsgroup. This retention period is defined as the period over which no new items were posted to the newsgroup, after which the newsgroup is automatically deleted. If not specified, the system default value (normally 60 days) is used.

If the value /HOLD=* is used, the newsgroup will NOT be automatically deleted.

6.15.2 /ITEMHOLD=n

This qualifier specifies the default retention period (in days) of the items within this newsgroup. All news items will remain on the system for the specified period, after which they will be automatically deleted. If not specified the system default value (normally 30 days) is used. Note that the SET ITEM command can be used to alter the retention period for an individual news item.

If the value /ITEMHOLD=* is used, NO items within the newsgroup will be automatically deleted.

6.15.3 /MODERATOR=address

This qualifier creates a moderated newsgroup. No user may post news items directly into a moderated newsgroup (unless the user is the actual moderator of the newsgroup). Any items posted into such newsgroups are redirected into the mail systems, and are mailed directly to the moderator for approval.

The moderator address should be specified in the internet format (user@node.site.domain) - NOT Decnet format (the translation from internet addresses to the equivalent Decnet addresses is handled internally by NEWS).

6.15.4 /LOCAL

Newsgroups are defined by default to be Network newsgroups - A network newsgroup is created in response to a network-wide ”newgroup” control message, and is deleted in response to a network-wide ”rmgroup” or ”checkgroups” message.

Newsgroups which are local to a this site, or local to a domain within the wider network are created with the /LOCAL qualifier. This ensures that the newsgroup will not be deleted by the network control messages.

6.15.5 /RESTRICT

This qualifier creates a restricted newsgroup. A restricted newsgroup is one where access to the newsgroup is restricted to a list of members. The membership of the newsgroup is determined by the ADD MEMBER, DELETE MEMBER and MODIFY MEMBER commands.

6.15.6 /SERVER=node

Format:

    CREATE NEWSGROUP name/SERVER=node

Newsgroups may be linked to a remote NEWSSERVER via NNTP. The presence of a remote server for a newsgroup implies that the local site need not keep the text files of all items within a newsgroup on local storage – when a user on the local system reads an item, the remote server passes a copy of the text on demand.

This configuration can be used to reduce the overall storage demands for NEWS if a number of NNTP-connected systems are all running NEWS. One system may be nominated as the server for a number of client nodes, or a number of systems can be used to serve subsets of the total newsgroup set to the clients (i.e. a node may be configured as a client for some newsgroups, and a server for other newsgroups).

This qualifier requires a node name as a value – this name is the decnet node name of the remote server.

6.15.7 /PROTOCOL=proto

Specifies a protocol to be used to access the NNTP SERVER node. The values for this qualifier are ”DECNET”, ”CMUTCP” or ”WINTCP”. They indicate the transport protocol used to carry the NNTP traffic between this client and the server node.

6.15.8 /KEEPREQUEST

The /KEEPREQUEST qualifier is used in conjunction with the /SERVER qualifier. /KEEP-REQUEST is the default value.

If a newsitem is served from a remote system, and a local user requests to view the contents of that item, the server system copies the text of the item into the local user’s scratch area. By default, a copy of this text is also placed in the local NEWS database, so that further requests to read the item do not need to access the remote server.

This retention of the local copy can be inhibited by using the /NOKEEPREQUEST qualifier – in such a case all requests to view the item will generate remote server requests.

6.15.9 /FILEHOLD=n

The /FILEHOLD qualifier is used in conjunction with the /SERVER qualifier. /FILEHOLD=2 is the default value.

If a newsgroup is served from a remote system, then it is not necessary to keep local copies of the text of news items for the retention period of the news item record - as the server node will retain all items for the full period.

By default, the local text is deleted after 2 days, and any subsequent requests to view the text of the item will result in a remote server request for the item text (if /KEEPREQUEST is also enabled then the copy is again retained on the local system for a further 2 days, and so on). This text retention period can be set to n days using the /FILEHOLD qualifier.

6.15.10 /INQUIRE

This command is used in the context of a client NEWS system adding newsgroups into the client database to match those held by the client’s server system. The command allows the client system to import all, or a subset, of the newsgroups held on a news server node. The command requires the /SERVE and /PROTOCOL qualifiers as well (and if they are not specified, then the user will be prompted to supply the values).

The action of the command is to request the full list of newsgroups on the target server node. All newsgroups which do not already exist on the local node are then created, subject to confirmation by the user for each new newsgroup. The attributes of the created newsgroups are set to the server and protocol values as given in the command.

The newsgroup list specified in the command line is the pattern to match when looking for new newsgroups. The following examples illustrate the use of pattern specification.

To import all newsgroups defined on the node ”unixnode”, using the TCP transport, the command would be:

    CREATE NEWSGROUP */SERVE=unixnode/PROTO=TCP/INQUIRE

To import all newsgroups within the tree ”anu.*” from the node ”vmsnode”, using DECnet:

    CREATE NEWSGROUP anu.*/SERVE=vmsnode/PROTO=DECNET/INQUIRE

/INQUIRE also supports an additional qualifier, ”/SINCE=date”, which restricts the search on the remote system to those newsgroups created since the specified date.

6.15.11 /SINCE=date-timne

This command is used in conjunction with the /INQUIRE qualifier to restrict the search for new newsgroups on the server to those created since the specified time.

6.16 DEFINE/KEY

Format:

    DEFINE/KEY key-name equivalence-string

This command associates an equivalence string and a set of attributes with a key on the terminal keyboard.

The key-name parameter specifies the name of the key that you are defining. The following table gives the set of key names in column one. The remaining three columns indicate the key designations on the keyboards for the three different types of terminals that allow key definitions.

NOTE: that the arrow keys and the PREV SCREEN and NEXT SCREEN keys should not be re-defined from their default defined values.

Key-name       LK201       VT100-type      VT52
PF1            PF1         PF1             [blue]
PF2            PF2         PF2             [red]
PF3            PF3         PF3             [gray]
PF4            PF4         PF4             -
KP0,KP1...KP9  0,1...9     0,1...9         0,1...9
PERIOD         .           .               .
COMMA          ,           ,               n/a
MINUS          -           -               n/a
ENTER          Enter       ENTER           ENTER
E1             Find        -               -
E2             Insert Here -               -
E3             Remove      -               -
E4             Select      -               -
HELP           Help        -               -
DO             Do          -               -
F17...F20      F17...F20   -               -

The equivalence-string specifies the string which is to be processed when you press the key. If the string contains any spaces, enclose the equivalence string in quotation marks.

6.16.1 /ECHO

Determines whether or not the equivalence string is displayed on your screen after the key has been pressed. The default is /ECHO. You cannot use /NOECHO with the /NOTERMINATE qualifier.

6.16.2 /PROTECT

Determines whether or not the key definition can be changed or deleted once the definition is performed. /NOPROTECT is the default.

6.16.3 /IF_STATE=(state-name,...)

Specifies a list of one or more states, one of which must be in effect for the key definition to be in effect. If you omit the /IF_STATE qualifier or use /NOIF_STATE, the current state is used. The state name is an alphanumeric string. States are established with the /SET_STATE qualifier or the SET KEY command. If you specify only one state name, you can omit the parentheses. By including several state names, you can define a key to have the same function in all the specified states.

6.16.4 /LOCK_STATE

Specifies that the state set by the /SET_STATE qualifier remain in effect until explicitly changed. If you use the /NOLOCK_STATE qualifier, the state set by /SET_STATE is in effect only for the next defineable key that you press or for the next read terminating character that you type.

The default is /NOLOCK_STATE. The /LOCK_STATE qualifier can only be specified with the /SET_STATE qualifier.

6.16.5 /SET_STATE=state-name

Causes the specified state-name to be set when the key is pressed. The state name can be any alphanumeric string.

If you omit the SET_STATE qualifier or use /NOSET_STATE, the current state that was locked remains in effect. If you have not included this qualifier with a key definition, you can use the SET KEY command to change the current state.

6.16.6 /TERMINATE

Specifies whether or not the current equivalence string is to be terminated (that is, processed) when the key is pressed. The default is /NOTERMINATE, which allows you to press other keys before the equivalence string is processed. Pressing RETURN has the same effect as using /TERMINATE.

The /NOTERMINATE qualifier allows you to create key definitions that insert text into command lines, after prompts, or into other text that you are typing.

6.17 DELETE

DELETE is a NEWS MANAGER command.

DELETE is used to delete newsgroups and/or newsitems. If the screen is currently displaying the newsgroup directory, the DELETE command deletes the current newsgroup. If the newsitem directory is displayed on the screen, DELETE removes the current newsitem.

6.18 DELETE ENTRY

Format:

    DELETE ENTRY newsgroup

DELETE ENTRY is a user command, used to change the membership of a user class. The newsgroup parameter is the name of the newsgroup to remove. This newsgroup is removed from the current class (or the class specified by the /CLASS qualifier. To delete an entire class use the value ”*” for the newsgroup name.

6.18.1 /CLASS=class-name

specifies the class from which the newsgroup is to be removed. To remove a newsgroup from all classes use the value ”*” as the class-name value.

6.19 DELETE NEWSGROUPS

DELETE NEWSGROUPS is a NEWS MANAGEMENT command

Format:

    DELETE NEWSGROUPS newsgroup-list

This command deletes newsgroups from the local NEWS database. The command may only be used by the local NEWS manager. Any newsitems held within the specified newsgroups are also deleted.

DELETE GROUPS may also be used to delete newsgroups.

If no newsgroup names are specified, then the current newsgroup is deleted. Otherwise the newsgroups are specified as a list of names, separated by commas. The ’*’ wildcard character may also be used to specify a set of newsgroups.

6.19.1 /CONFIRM

This qualifier directs NEWS to prompt for confirmation for each newsgroup to delete. This value is on by default, and is removed by the ”/NOCONFIRM” qualifier.

6.20 DELETE NEWSITEMS

DELETE NEWSITEMS is a NEWS MANAGEMENT command

Format:

    DELETE NEWSITEMS newsitem-list
    DELETE NEWSITEM/ID="message-identifier"

This command deletes newsitems from the current newsgroup from the local NEWS database. The command may be used by the local NEWS manager, or by the local moderator of a newsgroup.

DELETE ITEMS and DELETE NOTE may also be used to delete newsitems.

By default the current newsitem is deleted, or a list of note numbers may be specified as a command parameter. The format of this list is a list of numbers, ranges (start-end), and the keywords ”first”, ”last”, ”all”, ”.” (current item), ”*” (all items).

For example, to delete items 3 through to 10, 20 and 22 to 25 the command is: DELETE NEWSITEMS 3-10,20,22-25. To delete all items in a newsgroup the command is: DELETE NEWSITEMS *.

6.20.1 /CONFIRM

This qualifier directs NEWS to prompt for confirmation for each newsgroup to delete. This value is on by default, and is removed by the ”/NOCONFIRM” qualifier.

6.20.2 /IDENTIFIER=”message-id”

The /IDENTIFIER qualifier is used to select a news item with a specified message identifier string. The format is ”/IDENTIFIER=message-id”. Note that the identifier value includes the ’<’ and ’>’ characters. If /IDENTIFIER is used, this excludes the use of a set of news item numbers in the command.

6.21 DELETE MARKER

Format:

    DELETE MARKER [tag]

DELETE MARKER removes all marks of the specified tag from the mark list. If no tag value is specified then ALL marks are cleared. Otherwise only the specified tag value is cleared. The wildcard character (’*’) can be used in the tag field to clear a set of related tags. See Also CLEAR MARK, MARK, UNMARK

6.22 DELETE MEMBER

DELETE MEMBER is a NEWS MANAGEMENT command. Format:

    DELETE MEMBER username

This command deletes the specified username entry from the access list of the current newsgroup.

6.23 DEREGISTER

Format:

    DEREGISTER newsgroup
    DEREGISTER/ALL

This command removes your registration from a newsgroup. The newsgroup may either be specified as a parameter to the command, or, if not specified, the current newsgroup is deregistered. The parameter to the command is the name of the newsgroup to deregister. If no newsgroup name is specified, the current newsgroup is deregistered. See Also REGISTER, DIR/REGISTER, DIR/NEW, SELECT/NEW

6.23.1 /ALL

The /ALL qualifier removes registration tags from all newsgroups.

6.24 DIRECTORY

In screen mode this command is used to move between newsgroup and newsitem displays, and also to select some subset of the full newsgroup set to display on the screen.

In line mode, this command, by default lists all the news items in the currently selected newsgroup, on a page by page basis, displaying the item number, creation date and item title.

6.24.1 /CLASS=class-name

Sets the current class to that specified by this qualifier. All other directory qualifiers are prefixed by the condition that the newsgroups must be members of the current class. To remove a class selection condition use the class name ”*”.

6.24.2 /ALL

In screen mode this command resets the newsgroup directory to display all newsgroups held on the local system.

In line mode these newsgroups are listed page by page.

6.24.3 /NEW

In screen mode this command sets the newsgroup directory screen to only display those newsgroups in which you are registered, and there are unread newsitems within those newsgroups.

In line mode these newsgroups are listed page by page.

6.24.4 /REGISTER

In screen mode this command sets the newsgroup directory screen to only display those newsgroups in which you are registered.

In line mode these newsgroups are listed page by page.

6.24.5 /SINCE=date

In screen mode this command sets the newsgroup directory screen to only display those newsgroups in which new items have been entered since the date specified. The date is given in VMS time format: e.g. DIR/SINCE=22-FEB.

In line mode these newsgroups are listed page by page.

6.24.6 /ITEMS

This opens the ’current’ newsgroup and displays the newsitem for that newsgroup. This command is equivalent to the basic ’SELECT’ command.

6.24.7 /NEWSGROUPS

This command moves the context of NEWS to the newsgroup directory screen. This command is equivalent to the ’NEWSGROUP’ command.

6.24.8 /RESET

This clears the internal memory copy of the NEWS database, and reads a new copy of the database from the master index files into memory.

This command is not normally used – it is relevant only in those situations where a background process is adding or deleting news items to the database, and the user encounters ’file not found’ errors when attempting to access item files.

6.25 DOWN

Format

    DOWN [number]

In screen mode this moves the current pointer down one line. The command has an optional numeric parameter, the number of lines to move down. If no number is specified, 1 is the default value. This command is bound to the down arrow key on the terminal. The command ’DOWN 18’ is bound to the ’next screen’ key in the terminal.

6.26 EDIT

Format:

    EDIT filename

Invokes the Editor to edit a VMS file. On exit from the editor, NEWS is resumed.

6.27 EXIT

EXIT from NEWS. Ctrl-Z is also interpreted as EXIT.

On exit from NEWS a new register file is written out to SYS$LOGIN. This file, ”newsrc”, contains the list of registered newsgroups, read/unread newsitems, the marked item list and the kill filter lines.

The context file is used to determine which news items are unread, and is also used to re-establish context on the next invocation of NEWS – The next invocation of NEWS will restore the screen displays to the state at which the ’EXIT’ command was entered. On exit any pending print requests are released into the print queues.

See Also QUIT

6.28 EXTRACT

Format:

    EXTRACT [output-file [item-range]]

EXTRACT extracts a news item into a text file. The output-file parameter specifies the name of the file to save to, and the item-range specifies the range of item numbers to save.

6.28.1 /ALL

Specifies that all items in the current newsgroup are to be extracted to the specified file.

6.28.2 /APPEND

Appends the items to the specified file. If the file does not exist, it is created.

6.28.3 /AUTHOR=mailaddress

Specifies that only those items (from the specified item-range) posted from the specified mail address are to be saved. This qualifier is a synonym for /FROM.

6.28.4 /BEFORE=date-time

Specifies that only those items (selected from the specified item-range) postmarked before the specified date are to be saved.

6.28.5 /FROM=mailaddress

Specifies that only those items (from the specified item-range) posted from the specified mail address are to be saved. This qualifier is a synonym for /AUTHOR.

6.28.6 /HEADER

Format:

    EXTRACT/HEADER     [default]

    EXTRACT/NOHEADER

Specifies whether or not to include the full item headers in the saved text.

6.28.7 /MARKER

Format:

    EXTRACT/MARKER[=tag]
    EXTRACT/MARKER=(tag-list)

Specifies that only those items marked with one of the specified tag values are to be saved.

6.28.8 /SEEN

Marks the extracted item as seen.

6.28.9 /SINCE=date-time

Specifies that only those items (selected from the specified item-range) postmarked after the specified date are to be extracted.

6.28.10 /SUBJECT=”string”

Extract only those items (selected from the specified item-range) which contain the specified string within the item subject line. This qualifier is a synonym for /TITLE.

6.28.11 /TITLE=”string”

Extract only those items (selected from the specified item-range) which contain the specified string within the item subject line. This qualifier is a synonym for /SUBJECT.

6.28.12 /UNREAD

Extract only those items (selected from the specified item-range) which have not been read. This qualifier is a synonym for /UNSEEN.

6.28.13 /UNSEEN

Extract only those items (selected from the specified item-range) which have not been read. This qualifier is a synonym for /UNSEEN.

6.29 FOLLOWUP

FOLLOWUP is used to post a reply to a news item back into NEWS. The followup item will reference the news item you are following up as the parent of the new posting.

The editor is invoked to form the new posting, with the edit buffer preloaded with the text of the original item. The default editor is the TPU EDT emulator, but this is user definable - see the section on customisation to specify a different editor to use.

NEWS will fill in all defaulting news headers, and will prompt for all header items which are non-defaulting (the subject line and the newsgroups to post the item into).

If the newsgroups for the posting are moderated newsgroups, the posting is automatically redirected as mail to the moderator address, and NOT posted back to NEWS.

See Also REPLY, POST

6.29.1 /HEADERS

Followup will normally prompt for the subject of the posting, and a list of newsgroups to post the item into. All other header lines of the news item text will be filled with null values.

The /HEADERS qualifier will force prompting for optional components of the item header. These fields are:

Summary: A one line summary of the contents of the posting.

Reply-To: The (internet) mail address for REPLYing to the item - used if replies are to be sent to a mail address OTHER than yourself (as the originator of the item.

Distribution: A (comma separated) list of keywords designed to limit the distribution of a news item to a subset of the network. These keywords are site specific, but commonly include the keywords ”local” (do not forward to the net) and ”world” (forward as far as possible through the net).

Followup-To: The list of newsgroups where followup postings are to be sent. This is used to move a discussion to another newsgroup.

Keywords: A few relevant, pithy keywords which should indicate to other readers if the item is of interest.

6.30 FORWARD

Invokes VMS MAIL, to send a copy of the current news item via mail to another user. FORWARD and MAIL are equivalent. See Also MAIL

6.30.1 /TO=user

Format:

 FORWARD/TO=user
 FORWARD/TO=(user,user)

Specifies the VMS Mail address of the users to receive the mail (This is one of the anomalies of NEWS that will be addressed in a future version - the addresses used here should be internetstyle addresses - not VMS Mail addresses -as is currently the case). This may be a user name, or a list of names, and may also include a distribution list file specification. If this qualifier is not specified, the user is prompted for the /TO value.

6.30.2 /SUBJECT=”subject”

Format:

 FORWARD/SUBJECT="text string"

Specifies the subject of the message for the heading. If not specified, the user is prompted for the /SUBJECT value.

6.30.3 /NOEDIT

By default, FORWARD invokes the editor to allow the item text to be edited before mailing the item. This qualifier sends the news text to the mailer without modification. The default editor is the TPU EDT emulator, but this is user definable - see the chapter on customisation of NEWS.

6.30.4 /SELF

Sends a copy of the message back to yourself, as well as those on the /TO list.

6.30.5 /HEADERS

By default the news item header lines are stripped off the item text before it is mailed. This qualifier loads the item header lines into the mail buffer as well as the body of the text.

6.31 HELP

This command invokes the help processor. The command can be used with parameters to specify the command on which you want the help displayed.

6.32 KILL

The KILL facility allows the user to filter out all messages of a particular class - the READ/NEW command will mark such filtered items are having been read without displaying the text of the item to the user, and search forward for the next unread item. This enables the user to skip messages whose subject, or sender, are of no particular interest.

The KILL filter works on four possible attributes of an item for a match with the filters: the newsgroup name, the subject line, the sender’s address and any other single item header line. Of these attributes, only the newsgroup name and a minimum of one of the remaining three filters need be specified (If more than one item attributed is specified then the logical AND of the filters is the result of the filter). Thus it is possible to filter all postings by a particular sender in all newsgroups, or filter a conversation stream from a particular newsgroup, or filter all postings from a particular sender within a specified conversation stream.

The KILL command adds entries to the kill filter list. Removing kill filters is performed by the CLEAR KILL command, and existing kill filters can be edited using the MODIFY KILL command.

The KILL command has three qualifiers: /SUBJECT, /FROM and /HEADER. The /SUBJECT qualifier directs NEWS to add the current item’s subject line to the kill filter set, and the /FROM directs NEWS to add the current item’s ”From:” line to the kill filter set. The /HEADER qualifier takes a text pattern as a value, and matches the pattern against a header line. This header line value is taken as the filter value.

The command will then prompt for the set of newsgroups to which this filter will be applied. The default value is the current newsgroup however more newsgroups may be specified using wildcard notation (’*’) or a list of newsgroups.

See Also SHOW KILL, MODIFY KILL, CLEAR KILL

6.32.1 /FROM

Add the current item’s From: address to the kill filter set.

6.32.2 /SUBJECT

Add the current item’s Subject line to the kill filter set.

6.32.3 /HEADER=”text”

The /HEADER qualifier takes a text pattern as a value, and matches the pattern against a header line. This header line value is taken as the filter value.

For example, to kill all items with the same set of keywords as the current item the command would be:

    KILL/HEADER="keywords:"

This command will read the current item, extract the keywords header line, and use this line as a kill filter.

6.33 MAIL

See the section on the FORWARD command: FORWARD and MAIL are equivalent.

6.34 MARK

Format:

    MARK [tag] [item-number]

MARK places a mark against the current newsitem. This item can be selected at a later stage by the READ/MARK or SELECT/MARK commands.

If no tag value is specified then the default tag, ”mark” is used. Any tag value may be specified - tag values are words (no space characters) of any (well less than 80) length.

If an item number is specified then the specified item is marked, rather than the default of the current item. See Also UNMARK, CLEAR MARK

6.35 MODIFY ENTRY

Format:

    MODIFY ENTRY newsgroup/CLASS=(class-list)

Alters the classes to which the newsgroup is assigned.

6.35.1 /CLASSES=(class-list)

Specifies a new list of classes to which the specified newsgroup is assigned.

6.36 MODIFY KILL

Format:

    MODIFY KILL tag

MODIFY kill allows existing kill filters to be edited. The kill filter with the nominated tag is displayed, then each of the four filter attributes (sender, subject, newsgroup and header line pattern) can be edited (using the command line editor functions). If all four attributes are cleared then the filter is deleted. See Also KILL, SHOW KILL, CLEAR KILL

6.37 MODIFY MEMBER

Format:

    MODIFY MEMBER username

This command allows a newsgroup moderator, or the local NEWS manager to modify an entry in the membership list of a restricted newsgroup.

6.37.1 /MAIL=”mail-address”

Sets or changes the e-mail address of the user.

6.37.2 /NAME=new-username

Modifies the username of the entry.

6.37.3 /PRIVILEGES=(new-priv-list)

Sets the user’s newsgroup-privileges to the specified values. The values are: MODERATE (grant local moderator privilege), WRITE (allow write access to a write-protected newsgroup) and NOWRITE (disallow write access).

6.38 NEWSGROUPS

In screen mode this pops the display stack back to the newsgroup directory screen from the newsitem or news text display screens. This command is a synonym of the ’DIR/NEWSGROUPS’ command.

6.39 NOSCREEN

This command turns off screen mode and enters line command mode. See Also SCREEN

6.40 OPEN

Format:

    OPEN [newsgroup]

OPEN changes newsgroup context to the nominated newsgroup. If no newsgroup is given, the command selects the ’current’ newsgroup. Otherwise the specified newsgroup is opened. For example, ”OPEN test” opens newsgroup ’test’. ”OPEN” and ”SELECT” are synonyms.

6.40.1 /NEW

This qualifier is used to select the next registered newsgroup which contains unseen news items. If the current newsgroup contains unread items, then no action is performed.

6.40.2 /MARK

Format:

    OPEN/MARK[=tag]

This command moves the current item pointer to the NEXT newsitem marked with the specified tag value. If no tag value is specified than the next item marked with any tag is used.

6.41 POST

Format:

    POST [filename]

POST is used to post a new news item into NEWS. By default, the editor is invoked to form the new posting. The default editor is TPU, but this is user definable - see the chapter on customisation to specify a different editor to use.

NEWS will fill in all defaulting news headers, and will prompt for all header items which are non-defaulting (the subject line and the newsgroups to post the item into).

If the newsgroups for the posting are moderated newsgroups, the posting is automatically redirected as mail to the moderator address, and NOT posted directly to NEWS.

If a filename is specified, the editor is loaded with the text of the specified file, otherwise the editor is invoked to create a new file.

See Also FOLLOWUP, REPLY, FORWARD

6.41.1 /EDIT

The default action is to call the editor to create the posting. If you specify a file as a parameter to the WRITE command, and also specify the qualifier /NOEDIT , no call is made to the editor.

6.41.2 /HEADERS

Post will normally prompt for the subject of the posting, and a list of newsgroups to post the item into. All other header lines of the news item text will be filled with null values.

The /HEADERS qualifier will force prompting for optional components of the item header. These fields are:

Summary: A one line summary of the contents of the posting.

Reply-To: The (internet) mail address for REPLYing to the item - used if replies are to be sent to a mail address OTHER than yourself (as the originator of the item.

Distribution: A (comma separated) list of keywords designed to limit the distribution of a news item to a subset of the network. These keywords are site specific, but commonly include the keywords ”local” (do not forward to the net) and ”world” (forward as far as possible through the net).

Followup-To: The list of newsgroups where followup postings are to be sent. This is used to move a discussion to another newsgroup.

Keywords: A few relevant, pithy keywords which should indicate to other readers if the item is of interest.

Expires: A date to indicate the end of the useful life of the information in the message. NEWS will reject expirations greater than 12 months as being too outlandish.

6.41.3 /NEWSGROUPS

Format:

 POST/NEWSGROUPS=newsgroup-list

This qualifier specifies the list of newsgroups to receive the posting. If this qualifier value is not specified, the user is prompted for the newsgroups. The list is a comma separated list of newsgroup names.

6.41.4 /SELF

If specified a copy of your posting is mailed to yourself via VMS MAIL.

6.41.5 /SUBJECT=subj

Format:

 POST/SUBJECT="text string"

Specifies the subject of the message for the heading. If not specified, the user is prompted for the /SUBJECT value.

6.42 PRINT

Format:

    PRINT [item-range]

PRINT submits a print job into the print queues.

Print appends the current news item to a temporary print file. As more items are PRINTed, they are appended to this print file.

By default the print job is only submitted at the end of the NEWS session, but this can be altered, by specifying the /NOWAIT qualifier, to an immediate print job submission.

If no item-range is specified the current newsitem is submitted.

There are three qualifier sets: qualifiers which affect the print mode, qualifiers which select the items and their print formats, and the standard VMS print job qualifiers.

Print mode is by default to print newsitem contents (item mode). The /DIRECTORY qualifier is used to select directory mode, where NEWS directory listings may be printed.

If the item mode is used then the default is to print the current newsitem. This behaviour may be altered by selecting a range of items, with selection criteria also applied by qualifiers.

6.42.1 /WAIT

Specifies whether the job is to be submitted at the end of the current NEWS session (/WAIT [default]) or printed immediately (/NOWAIT).

6.42.2 /DIRECTORY=option [mode]

This qualifier modifies the PRINT command to print a directory listing of all newsgroups, or of all newsitems within a newsgroup instead of printing a newsitem.

The options available with this qualifier are:

• NEWSGROUPS prints a directory listing of all newsgroups. ”GROUPS” is an alternative option value for this action.

• NEWSITEMS prints a directory listing of the current newsgroup. ”ITEMS” and ”NOTES” are alternative option values for this action.

6.42.3 /NEWSGROUPS=(group-list) [directory]

Specifies the newsgroups for which the directory mode is to apply. By default the current newsgroup is used.

6.42.4 /ALL [item]

Selects all news items in the current newsgroup for printing. If the newsgroup contains a large number of items this could seriously affect your disk quota (as well as risking the possibility of printer meltdown!).

6.42.5 /AUTHOR=mailaddress [item]

Specifies that only those items (from the specified item-range) posted from the specified mail address are to be printed. This qualifier is a synonym for /FROM.

6.42.6 /BEFORE=date-time [item]

Specifies that only those items (selected from the specified item-range) postmarked before the specified date are to be printed.

6.42.7 /CONFIRM [item]

As with the normal VMS /CONFIRM qualifier semantics, the user is required to confirm each newsitem to be printed.

6.42.8 /FROM=mailaddress [item]

Specifies that only those items (from the specified item-range) posted from the specified mail address are to be printed. This qualifier is a synonym for /AUTHOR.

6.42.9 /HEADER [item]

Format:

    PRINT/HEADER     [default]
    PRINT/NOHEADER

Specifies whether or not to include the full item headers in the printed text.

6.42.10 /MARKER [item]

Format:

    PRINT/MARKER[=tag]
    PRINT/MARKER=(tag-list)

Specifies that only those items marked with one of the specified tag values are to be saved.

6.42.11 /SEEN [item]

Marks the printed item as seen.

6.42.12 /SINCE=date-time [item]

Specifies that only those items (selected from the specified item-range) postmarked after the specified date are to be printed.

6.42.13 /SUBJECT=”string” [item]

Print only those items (selected from the specified item-range) which contain the specified string within the item subject line. This qualifier is a synonym for /TITLE.

6.42.14 /TITLE=”string” [item]

Print only those items (selected from the specified item-range) which contain the specified string within the item subject line. This qualifier is a synonym for /SUBJECT.

6.42.15 /UNREAD [item]

Print only those items (selected from the specified item-range) which have not been read. This qualifier is a synonym for /UNSEEN.

6.42.16 /UNSEEN [item]

Print only those items (selected from the specified item-range) which have not been read. This qualifier is a synonym for /UNSEEN.

6.42.17 /AFTER=date-time [print]

Specifies that the job will be held in the print queue until after the specified time.

6.42.18 /BURST=ALL—ONE [print]

Specifies whether a burst page is to be printed.

6.42.19 /CHARACTERISTICS=char-value [print]

Specifies the characteristics for the print job.

6.42.20 /COPIES=n [print]

Specifies the number of copies of each item to be printed.

6.42.21 /FEED [print]

Specifies whether formfeeds are to be inserted into the print stream by the printing software.

6.42.22 /FLAG=ALL—ONE [print]

Specifies whether an item flag page is to be printed.

6.42.23 /FORM=form [print]

Specifies the form to use for the print job.

6.42.24 /HOLD [print]

HOLD the request until manually released.

6.42.25 /JOBCOUNT=n [print]

Specifies the number of repeat prints of the job.

6.42.26 /LOWERCASE [print]

Specifies that the job is to be printer in dual case.

6.42.27 /NAME=job-name [print]

Specifies a job name for the submitted print job.

6.42.28 /NOTE=”job-note” [print]

Attaches a note to the submitted print job.

6.42.29 /NOTIFY [print]

Requests VMS to broadcast a message to you on print completion.

6.42.30 /OPERATOR=”operator request” [print]

Tags the print job with an operator request.

6.42.31 /PASSALL [print]

Specifies whether the job is to be printed in ”raw” mode.

6.42.32 /PARAMETERS=(param-list) [print]

Specifies job print parameters.

6.42.33 /PRIORITY=priority [print]

Specifies a print job priority. You need VMS privileges to specify a value higher than the system default.

6.42.34 /QUEUE=queue-name [print]

Specifies a print queue (other than the default queue) for printing.

6.42.35 /SETUP=setup-modules [print]

Specifies setup modules for the print job.

6.42.36 /TRAILER=ALL—ONE [print]

Specifies whether a trailer page is to be printed for the job.

6.43 QUIT

QUIT exits from news without updating the user’s news register file. Thus, no changes are made to the set of unread items, nor are any changes made to the set of marked news items (this is a more graceful means of Ctrl-C or Ctrl-Y!).

See Also EXIT

6.44 READ

This command reads the next item from the current selected newsgroup, and displays the item on the screen page-by-page.

The actions of READ can be modified by giving a Newsitem number, which will display that item.

6.44.1 READ newsitem_number

This parameter may be either the number of a news item, a ’.’ (to re-read the last read item from the start), or ’*’ to read the last item in the group.

6.44.2 /HEADER

Normally NEWS will suppress the display of all network routing headers. This qualifier produces a full display of the item contents including mail header lines.

6.44.3 /NEW

This command will display the next unread NEWS item from the set of registered newsgroups. This may cause an implicit SELECT of a different newsgroup to obtain the next unread item.

6.44.4 /EDITOR

This qualifier will use an editor in read-only mode to view the newsitem. This is useful when it is required to scroll up and down through the item text, or extract a part of the text into a file.

The default editor is the TPU editor, but this is user definable - see the chapter on customisation of NEWS.

6.44.5 /TPU

This qualifier is equivalent to the /EDITOR qualifier. This qualifier will use an editor in readonly mode to view the newsitem. This is useful when it is required to scroll up and down through the item text, or extract a part of the text into a file.

The default editor is the TPU editor, but this is user definable - see the chapter on customisation of NEWS.

6.44.6 /NEXT

When reading a news item this command causes the remainder of the text display to be skipped, and the next item to be displayed.

6.44.7 /LAST

When reading a news item this command causes the remainder of the text display to be skipped, and the previous item to be displayed.

6.44.8 /BACK

This qualifier is equivalent to the /LAST qualifier. When reading a news item this command causes the remainder of the text display to be skipped, and the previous item to be displayed.

6.44.9 /PARENT

NEWS items may reference a previous item that has been posted. In such a case the text of the item commonly begins with the line:

”In article <identifier>, username writes:”

When a news item references previous items, the header includes the identifier of the parent item. READ/PARENT will display the parent item referenced by the current item.

6.44.10 /IDENTIFIER=”id”

All NEWS items are identified by a unique message identification string. (This identifier can be displayed using the READ/HEADER command, and noting the ”Message-Id: ...” line).

An item to be displayed can be selected by providing the identifier of the item. e.g.:

    READ/IDENTIFIER="<243@csc.anu.oz>"

6.44.11 /ROT13

Some news items are posted using a basic encryption method (rot13). This qualifier uses the decryption algorithm to display the message contents.

6.44.12 /MARKER

Format:

    READ/MARKER[=tag]

This command displays the text of the NEXT newsitem marked with the specified tag value. If no tag value is specified than the next item marked with any tag value is used.

6.44.13 /FOLLOWUP

This command scans the current newsgroup for the next newsitem with a subject field which matches that of the current newsitem subject (within the current newsgroup). In this way the responses (if any) to the current newsitem can be viewed.

6.44.14 /PREV

NEWS maintains a history of the last 10 items read. READ/PREV pops the most recent item index off this history stack and displays the item. Repeated calls to READ/PREV will continue to pop the stack until the history stack is emptied.

6.44.15 /TOPIC

READ will display the oldest posted newsitem in the same conversation stream as the current item. The oldest item will be original posting in the conversation stream if the original item is still retained on the system.

6.45 REFRESH

In screen mode this command repaints the screen. This command is (by default) bound to the Ctrl-W key.

6.46 REGISTER

Format:

    REGISTER [newsgroup-name]

Adds newsgroups into the user’s register file. Registering into a newsgroup implies that NEWS will automatically keep track of which items in the newsgroup have been read, and the user can view only the unread news items with repeated READ/NEW commands.

The newsgroup-name parameter specifies the name of the newsgroup to register. If the parameter is not specified, the current newsgroup is added to the register file.

See Also DEREGISTER

6.46.1 /CONFIRM

Prompt for confirmation before registering each newsgroup. Default is /NOCONFIRM.

6.46.2 /ALL

This qualifier will register the user for ALL newsgroups on the system. On a system with a high quantity of network news this must be considered a brave act!

6.46.3 /BEFORE=date-time

Registers all newsgroups which were created before the specified time.

6.46.4 /LOCAL

Registers all newsgroups which are defined as LOCAL newsgroups.

6.46.5 /NETGROUPS

Registers all newsgroups which are defined as NETWORK newsgroups (i.e. the complement of /LOCAL).

6.46.6 /NEW

This qualifier will register the user for all newsgroups created on the system since the last time a REGISTER/ALL or REGISTER/NEW command was given. Thus, the user is automatically registered into all recent newsgroups, and can then elect to deregister from some of those newsgroups.

6.46.7 /PRIORITY=n

This qualifier assigns a read priority to the newsgroup. This is particularly useful if you are registered in a number of newsgroups, and would like to view new items (using READ/NEW) in a particular order of newsgroups. The value is a number (between 1 and 255), increasing with priority.

For example, if you have new items in net.a and net.b and the priority of net.b is greater then net.a, unread items in net.b will be displayed first (when using READ/NEW.

6.46.8 /RECORD

Record the current time as the most recent register time. This will alter the behaviour of the next REG/NEW command such that the /NEW qualifier will only select newsgroups which have been created after this command was executed.

/RECORD is the default with the /NEW qualifier - /NORECORD is the default with all other qualifiers.

6.46.9 /SINCE=date-time

Registers all newsgroups which were created since the specified time.

6.47 REPLY

REPLY posts a MAIL reply directly to the sender of a news item. This is used to respond directly to the poster of an item without the item appearing on the net.

News will pre-load the edit buffer with the text of the item to which the reply is being sent, then invoke the editor. The default editor is the TPU EDT emulator, but this is user definable - see the section on customisation of NEWS.

On a normal exit from the editor the item is posted via VMS Mail to the sender (The sender’s address is mapped into the equivalent VMS Mail address by site specific address transformation rules).

See Also FORWARD, FOLLOWUP

6.47.1 /HEADERS

By default the news item headers are stripped off the item text before it is mailed. This qualifier loads the item header lines into the mail buffer as well as the body of the text.

6.47.2 /SUBJECT=”mail-subject”

This specifies the subject header of the mail reply. If not specified, NEWS will prompt for a ”Subject:” field for the mail.

6.47.3 /SELF

This is a mail qualifier which will forward a copy of the outgoing mail item back to the sender.

6.48 SAVE

Format:

    SAVE [output-file [item-range]]

SAVE extracts a news item into a text file. The output-file parameter specifies the name of the file to save to, and the item-range specifies the range of item numbers to save.

6.48.1 /ALL

Specifies that all items in the current newsgroup are to be extracted to the specified file.

6.48.2 /APPEND

Appends the items to the specified file. If the file does not exist, it is created.

6.48.3 /AUTHOR=mailaddress

Specifies that only those items (from the specified item-range) posted from the specified mail address are to be saved. This qualifier is a synonym for /FROM.

6.48.4 /BEFORE=date-time

Specifies that only those items (selected from the specified item-range) postmarked before the specified date are to be saved.

6.48.5 /FROM=mailaddress

Specifies that only those items (from the specified item-range) posted from the specified mail address are to be saved. This qualifier is a synonym for /AUTHOR.

6.48.6 /HEADER

Format:

    SAVE/HEADER     [default]

    SAVE/NOHEADER

Specifies whether or not to include the full item headers in the saved text.

6.48.7 /MARKER

Format:

    SAVE/MARKER[=tag]
    SAVE/MARKER=(tag-list)

Specifies that only those items marked with one of the specified tag values are to be saved.

6.48.8 /SEEN

Marks the saved item as seen.

6.48.9 /SINCE=date-time

Specifies that only those items (selected from the specified item-range) postmarked after the specified date are to be saved.

6.48.10 /SUBJECT=”string”

Save only those items (selected from the specified item-range) which contain the specified string within the item subject line. This qualifier is a synonym for /TITLE.

6.48.11 /TITLE=”string”

Save only those items (selected from the specified item-range) which contain the specified string within the item subject line. This qualifier is a synonym for /SUBJECT.

6.48.12 /UNREAD

Save only those items (selected from the specified item-range) which have not been read. This qualifier is a synonym for /UNSEEN.

6.48.13 /UNSEEN

Save only those items (selected from the specified item-range) which have not been read. This qualifier is a synonym for /UNSEEN.

6.49 SEARCH

Format:

    SEARCH [target string]

This command directs NEWS to locate a newsitem which contains a specified text string. The next item which contains the text string is displayed on the screen. Repeated calls to SEARCH (without specifying a text string) will locate successive occurrences in other news items that are in the specified newsgroups.

If no target string is specified, the search is resumed using the previous target string.

The search is usually case-insensitive, but this may be reversed (made case sensitive) by specifying a mixed-case string (i.e. upper case only and lower case only target strings are caseinsensitive). To specify a mixed-case string you must enclose the string in ” characters (e.g. SEARCH ”String”).

As this involves examining the contents of all specified item files, this command may take a reasonably long time. The search may be interrupted by pressing the RETURN key. This will call up a NEWS input prompt, and the search may either be resumed (with another RETURN) or a new command may be given.

6.49.1 /NEWSGROUPS

Format:

    SEARCH/NEWSGROUPS=newsgroups

By default the search is limited to the current newsgroup. This can be modified by specifying the newsgroups which should be searched.

The wildcard character (’*’) may be used to include a set of newsgroups.

A number of newsgroups may be specified with the usual list notation (comma separated list).

For example:

    SEARCH/NEWSGROUPS=(aus.*,anu.*) "Wombats"

6.49.2 /EDIT

This qualifier will use an editor in read-only mode to view the newsitem. This is useful when it is required to scroll up and down through the item text, or extract a part of the text into a file.

The default editor is TPU, but this is user definable - see the section on customisation.

6.49.3 /HEADER

Normally NEWS will suppress the display of all network routing headers. This qualifier produces a full display of the item contents including mail header lines.

6.49.4 /NODISPLAY

If this qualifier is used, the screen directory will be positioned on the item which contains the target string, but item text will not be displayed.

6.50 SCREEN

This enters NEWS screen display mode (the default mode). This command is normally given sometime after a NOSCREEN command. See Also NOSCREEN

6.51 SELECT

Format:

    SELECT [newsgroup]

SELECT changes newsgroup context to the nominated newsgroup. If no newsgroup is given, the command selects the ’current’ newsgroup. Otherwise the specified newsgroup is opened. For example, ”SELECT test” opens newsgroup ’test’.

6.51.1 /NEW

This qualifier is used to select the next registered newsgroup which contains unseen news items. If the current newsgroup contains unread items, then no action is performed.

6.51.2 /MARK

Format:

    SELECT/MARK[=tag]

This command moves the current item pointer to the NEXT newsitem marked with the specified tag value. If no tag value is specified than the next item marked with any tag is used.

6.52 SEND

Format:

    SEND [file-name]

SEND creates and sends a VMS MAIL message from within NEWS.

SEND will, by default, prompt for the To: address and Subject line, and invoke the editor to create the MAIl message.

You can send a mail message to the author of the current news item you’re reading by using the /AUTHOR qualifier, and you can send a mail item to all members of a restricted newsgroup by using the /MEMBERS qualifier.

You can also pre-load the edit buffer with the text of the message you’re reading by using the /EXTRACT qualifier.

6.52.1 /AUTHOR

Directs SEND to address the mail to the address of the user who posted the message you are currently reading.

6.52.2 /EDIT

/EDIT is the default value of this qualifier, which directs SEND to call the editor to assemble the mail message. /NOEDIT bypasses this call to the editor.

6.52.3 /EXTRACT

Directs SEND to load the edit buffer with the text of the item you are currently reading.

6.52.4 /HEADERS

Used with /EXTRACT, this qualifier directs NEWS to load all message headers into the message buffer. Otherwise a subset of the message headers are included into the edit buffer.

6.52.5 /QUOTE

Format:

    SEND/EXTRACT/QUOTE="string"
    SEND/EXTRACT/NOQUOTE

Used with /EXTRACT, this qualifier is used to alter the quote character used to delimit the original item in the edit buffer. By default the original item is loaded using the ”> ” string at the start of each line 1 to indicate included text.

Using /QUOTE=”string”, an arbitrary string may be used as the quotation string. Using /NOQUOTE includes the item text verbatim.

6.52.6 /MEMBERS

If the current newsgroup is configured with a membership list then the mail message will be addressed to all members of this list.

6.52.7 /SELF

By default a copy of the mail message will NOT be sent to yourself. The /SELF qualifier overrides this default.

6.52.8 /SUBJECT

Format:

    SEND/SUBJECT="string"

SEND will normally prompt for the subject line of the mail message. The /SUBJECT qualifier is used to supply a subject string to the procedure.

6.52.9 /TO

SEND will normally prompt for the addresses to mail. These addresses may be specified using a command line parameter of the form: /TO=”address-list”

6.53 SET ACCESS

SET ACCESS is a NEWS MANAGEMENT command.

Format:

    SET ACCESS newsgroup

SET ACCESS allows the local news manager to define a VMS access control list to control read and write access to a newsgroup via the VMS ACL mechanisms.

NOTE: access control is managed more consistently by using the SET NEWSGROUP/RESTRICT=MEMBERS and then managing a membership list using the ADD/DELETE/MODIFY/SHOW MEMBER commands. The SET ACCESS command is not recommended as the best means of access control for NEWS.

SET ACCESS invokes the VMS ACL Editor to allow the local news manager to define or modify the access control list associated with the newsgroup. The access control list is placed on the matching directory in the NEWS data area.

6.54 SET BROADCAST_TRAPPING

Format:

    SET BROADCAST_TRAPPING  [default]
    SET BROADCAST_TRAPPING ON
    SET NOBROADCAST_TRAPPING
    SET BROADCAST_TRAPPING OFF

NEWS by default has broadcast trapping enabled. All broadcast messages to the terminal are trapped by NEWS, and displayed at the base of the screen, in order to avoid overwriting of parts of the screen.

This trapping may be disabled by specifying BROADCAST_TRAPPING OFF or NOBROAD-CAST_TRAPPING

6.55 SET CLASS

Format:

    SET CLASS class-name

This command modifies the current directory selection criteria to display those newsgroups which have been entered into the specified class. To turn off class selection criteria, use the class name ”*” as the specified class. See Also DIRECTORY/CLASS

6.56 SET DISPLAY

This command is used to set display functions of NEWS. With this release of NEWS only one display setting is modifiable: SCANSIZE.

6.56.1 SCANSIZE=n

Format:

    SET DISPLAY/SCANSIZE=n

SCANSIZE determines the number of lines to display of the first page of an item text. When using a low speed terminal, it is often helpful to only display the first few lines of the item to determine whether to view the remainder of the item text.

This parameter can also be specified as a logical name ”NEWS_SCANSIZE”.

6.57 SET HISTORY

SET HISTORY is a NEWS MANAGEMENT command.

Format:

    SET HISTORY
    SET HISTORY ON
    SET NOHISTORY
    SET HISTORY OFF

NEWS normally maintains a history file of recent message identifiers which have been held on the local system. This file is referred to whenever a new item is added into NEWS to ensure that the item has not already been held on this system, so as to avoid looping news items in the NEWS network.

This history function can be disabled by the SET NOHISTORY or SET HISTORY OFF commands. It will remain disabled until explicitly re-enabled by the NEWS manager.

6.58 SET ITEM

SET ITEM is a NEWS MANAGEMENT command. SET ITEM modifies the newsgroup retention attributes of news items. Format:

     SET ITEM item-num/HOLD=n

This requires the number of a newitem (or defaults to the current newsitem). Only one qualifier is associated with this command, /HOLD.

6.58.1 /HOLD=n

This command qualifier defines the retention period for news items The value of the qualifier is specified in days. If a 0 value is given, this indicates that a higher level (newsgroup based) default value should be used. If a value of ’*’ is given, this indicates that the retention is PERMANENT – the item will never be deleted by the SKIM command.

6.59 SET LINE_EDITING

Format:

    SET LINE_EDITING
    SET LINE_EDITING ON
    SET NOLINE_EDITING  [default]
    SET LINE_EDITING OFF

By default VMS NEWS disables the cursor keys line editing functions. The up and down arrow cursor keys are bound to the UP and DOWN commands.

VMS line editing may be enabled with the SET LINE_EDITING. In this case the keypad key 5 is bound to the UP function, and keypad 2 is bound to the DOWN function, and the arrow cursor keys perform the normal line editing functions.

6.60 SET MANAGER

SET MANAGER is a NEWS MANAGEMENT command.

Format:

    SET MANAGER     [default]
    SET MANAGER ON
    SET NOMANAGER
    SET MANAGER OFF

A user who holds the NEWs_MGR_ID identifier is identified to NEWS as a local news manager, and can execute all the management commands.

These manager privileges can be disabled by the SET NOMANAGER command, and enabled at a later stage by the SET MANAGER command. Note that on startup manager privileges ARE enabled.

6.61 SET [NO]MODERATOR

Format:

    SET MODERATOR
    SET MODERATOR ON
    SET NOMODERATOR
    SET MODERATOR OFF

NEWS allows two types of moderation of a newsgroup: The USENET model of moderation is one where all postings into a moderated newsgroup are intercepted and mailed directly to the e-mail address of the moderator of the newsgroup. The moderator’s responsibility is to approve the posting and place it into the newsgroup with an approval stamp. A USENET moderator is identified by a match of the user’s e-mail address with the list of moderator addresses. Such a moderator is more accurately called an Approval Moderator.

VMS NEWS also has a second level of moderation of a newsgroup, which effectively allows a subset of the NEWS Manager’s privileges to be given to nominated users, effective for a particular newsgroup. Such a Local Moderator can control the membership list for a restricted newsgroup by adding, altering and removing member records, and cancelling messages that were posted from the local system into the controlled newsgroup.

Users who have such local moderator privileges do not have such privileges enabled at the start of the session: the privileges are enabled by the SET MODERATOR command for all locally moderated newsgroups. A user who is nominated as a Approval Moderator is automatically configured as a VMS NEWS local moderator if the Approval Moderator’s mail address is located on the local node.

6.62 SET NEWSGROUP

SET NEWSGROUP is a NEWS MANAGEMENT command.

SET NEWSGROUP modifies the control attributes of a newsgroup. control attributes include the lifetime of items on the local system, newsgroup moderator addresses, and whether the newsgroup is served from a remote NNTP server.

Newsgroups are automatically deleted from the NEWS system if the last new item in the newsgroup was posted to the group more than n days previously. There are three levels of default for this value: A global compile-time constant (defined in the file NEWSSITE.H, normally 60 days), which may be overridden by a dynamic global value (SET NEWSGROUP default/HOLD=n days), which may be overridden for individual newsgroups by a dynamic newsgroup specific value (SET NEWSGROUP name/HOLD=n days). The value is specified in units of days, but there is an additional value, ’*’ which specifies that the newsgroup is not to be automatically deleted.

Newsitems records are automatically deleted once they have remained on the system for a set number of days. This retention period has fours levels of default: A global compile-time constant (defined in the file NEWSSITE.H, normally 30 days), which may be overridden by a dynamic global value (SET NEWSGROUP */ITEMHOLD=n days), which may be overridden for specific newsgroups by a dynamic newsgroup default value (SET NEWSGROUP name/ITEMHOLD=n days), which may be overridden for individual news items (SET ITEM number/HOLD=n days). The value is specified in units of days, but there is an additional value, ’*’ which specifies that the newsgroup is not to be automatically deleted.

Format:

     SET NEWSGROUP/qualifiers
     SET NEWSGROUP newsgroup-name/qualifiers
     SET NEWSGROUP default/quals

If no newsgroup name is given with the command the command assumes the name of the ’current’ newsgroup. If the newsgroup is specified as ’default’ then the attributes referred to by this command are the system defaults.

The qualifiers accepted for the global default (’default’) are /ITEMHOLD, /HOLD, /KEEP-REQUEST and /FILEHOLD and in the case where a particular newsgroup is named, /MOD-ERATOR and /SERVER.

6.62.1 /HOLD=n

This command qualifier defines the retention period for newsgroups. The value of the qualifier is specified in days.

If a 0 value is given, this indicates that a higher level default value should be used. If a value of ’*’ is given, this indicates that the retention is PERMANENT – the newsgroup will never be automatically deleted by the SKIM command.

6.62.2 /ITEMHOLD=n

This command qualifier defines the default retention period for news items within a newsgroup. The value of the qualifier is specified in days.

If a 0 value is given, this indicates that a higher level default value should be used. If a value of ’*’ is given as the value for days, this indicates that the default newsitem retention is PERMANENT.

6.62.3 /MODERATOR

Format:

    SET NEWSGROUP/MODERATOR=address
    SET NEWSGROUP/MODERATOR
    SET NEWSGROUP/NOMODERATOR

This command sets the newsgroup as a moderated newsgroup, or clears the moderator condition.

Only the moderator of a moderated newsgroup may post items into that newsgroup – all other users who attempt to post to a moderated newsgroup will have the news posting intercepted and re-directed as e-mail to the moderator address.

The value of the /MODERATOR is an internet-style address of the moderator. In the case where no address is given, backbone addressing is assumed.

The /NOMODERATOR is used to clear the moderator condition.

6.62.4 /LOCAL

Format:

    SET NEWSGROUP name/LOCAL
    SET NEWSGROUP name/NOLOCAL

Newsgroups are defined by default to be Network newsgroups - A network newsgroup is created in response to a network-wide ”newgroup” control message, and is deleted in response to a network-wide ”rmgroup” or ”checkgroups” message.

Newsgroups which are local to a this site, or local to a domain within the wider network are created with the /LOCAL qualifier. This ensures that the newsgroup will not be deleted by such network control messages.

6.62.5 /RESTRICT

Format:

    SET NEWSGROUP name/RESTRICT=MEMBERS

Sets the newsgroup as a restricted access newsgroup. Making the newsgroup restricted limits the access to the newsgroup to those users who are entered as ”members” of the newsgroup (see ADD MEMBER). Users who are not listed as members of the newsgroup cannot read or write into this newsgroup.

6.62.6 /WRITE

By default all newsgroups are writeable by all users (there is a qualification to this assertion - a set of newsgroups can be setup as read-only by use of the NEWS_NETPOST/NEWS_NONETPOST identifiers - the newsgroup set refers to the /NOLOCAL newsgroups).

A newsgroup can be setup as a read-only newsgroup by the command:

    SET NEWSGROUP name/NOWRITE

Once a newsgroup has been set as /NOWRITE only those users with write permission can post to this newsgroup (see ADD MEMBER to grant write permission to a username).

6.62.7 /SERVER=node

Format:

    SET NEWSGROUP name/SERVER=node
    SET NEWSGROUP name/NOSERVER

Newsgroups may be linked to a remote NEWSSERVER via DECNET. The presence of a remote server for a newsgroup implies that the local site need not keep the text files of all items within a newsgroup on local storage – when a user on the local system reads an item, the remote server passes a copy of the text on demand.

This configuration can be used to reduce the overall storage demands for NEWS if a number of decnet-connected systems are all running NEWS. One system may be nominated as the server for a number of client nodes, or a number of systems can be used to serve subsets of the total newsgroup set to the clients (i.e. a node may be configured as a client for some newsgroups, and a server for other newsgroups).

This qualifier requires a node name as a value – this name is the decnet node name of the remote server.

6.62.8 /KEEPREQUEST

The /KEEPREQUEST qualifier is used in conjunction with the /SERVER qualifier. /KEEP-REQUEST is the default value.

If a newsitem is served from a remote system, and a local user requests to view the contents of that item, the server system copies the text of the item into the local user’s scratch area. By default, a copy of this text is also placed in the local NEWS database, so that further requests to read the item do not need to access the remote server.

This retention of the local copy can be inhibited by using the /NOKEEPREQUEST with the SET command – in such a case all requests to view the item will generate remote server requests.

6.62.9 /FILEHOLD=n

The /FILEHOLD qualifier is used in conjunction with the /SERVER qualifier. /FILEHOLD=2 is the default value.

If a newsgroup is served from a remote system, then it is not necessary to keep local copies of the text of news items for the retention period of the news item record.

By default, the local text is deleted after 2 days, and any subsequent requests to view the text of the item will result in a remote server request for the item text (if /KEEPREQUEST is also enabled then the copy is again retained on the local system for a further 2 days, and so on). This text retention period can be set to n days using the /FILEHOLD qualifier.

6.62.10 /PROTOCOL=p

The /PROTOCOL qualifier is used in conjunction with the /SERVER qualifier. Valid values for this qualifier are ”DECNET” or ”TCP”. The default value is ”DECNET”.

If a newsgroup is served from a remote system, then either DECNET or TCP may be used to access the remote system. This qualifier sets the protocol to use when performing the remote access.

6.63 SET PROFILE

Format:

    SET PROFILE

SET PROFILE is used to define a usage profile for NEWS. The profile may contain the specification of the editor to use with NEWS, the default action of the OPEN command, the class to be selected on NEWS startup, the personal name to use with all postings, a set of qualifiers to use with all PRINT commands, and a startup DIRECTORY setting.

Many of these settings can also be defined using logical names (see the chapter on customization of NEWS). The PROFILE settings take precedence over any such logical name settings.

6.63.1 /AUTOMATIC

Format:

    SET PROFILE/AUTOMATIC[={DIRECTORY | UNSEEN}]

This controls the default action of the OPEN and SELECT commands. by default the OPEN (and SELECT) command displays the directory of the opened newsgroup. By specifying UN-SEEN, the first unseen item will be displayed when the OPEN or SELECT command is executed.

6.63.2 /CLASS

Format:

    SET PROFILE/CLASS=class-name
    SET PROFILE/NOCLASS

By default NEWS will startup with the same selection conditions as were used when NEWS last exited. You can specify a class to select on startup by using the /CLASS qualifier.

The values of the classname are either a valid class name, or the string ”MAIN”. MAIN indicates that no class selection should be used, and all newsgroups should be included in the directory selection conditions.

The default behaviour can be restored by using the /NOCLASS parameter.

6.63.3 /DIRECTORY

Format:

    SET PROFILE/DIRECTORY=directory-switch
    SET PROFILE/NODIRECTORY

NEWS will normally retain the directory context across invocations of the program (i.e. If you had executed DIR/NEW before exiting NEWS, then when you next run NEWS, the initial newsgroup directory state will be restored to DIR/NEW). It is possible to specify a constant initial directory state by defining a value for this qualifier as one of the strings:

 ALL  NEW  REGISTER  SINCE  SINCE=vms_date

If a PROFILE value is specified then this value will override that of the logical name NEWS_DEFAULT_DIRECTORY.

6.63.4 /EDITOR

Format:

    SET PROFILE/EDITOR=editor
    SET PROFILE/NOEDITOR

This qualifier is used to link an editor to NEWS. The editor is used to generate new postings, and to edit forwarded mail, and (optionally) in readonly mode, to scan the item contents (READ/EDITOR).

By default NEWS links the TPU editor to NEWS, using the value of the logical name TPUSECINI (or TPU$SECTION, depending on the version of VMS used) as the section to use.

The value of this qualifier may be one of:

• EVE (or TPU) to use the TPU emulator with either the EVE section, or the section as defined by the appropriate TPU logical name.

• EDT to use the callable EDT interface

• A file specification (including the ”@” symbol) which specifies a spawned command procedure to use as the editor for NEWS.

6.63.5 /PERSONAL_NAME

Format:

    SET PROFILE/PERSONAL_NAME="personal-name-string"
    SET PROFILE/NOPERSONAL_NAME

Sets the personal name string. This name is appended to your network address in all NEWS postings.

If not specified as a PROFILE entry, the logical name NEWS_PERSONAL_NAME is used (if defined), and if there is no logical name your VMS MAIL profile is examined to extract your MAIL personal name.

6.63.6 /PRINT

Format:

    SET PROFILE/PRINT="print-qualifier-string"
    SET PROFILE/NOPRINT

Sets the default print qualifiers for all NEWS PRINT commands. The qualifiers used may be any valid NEWS print qualifier string (See the section describing the PRINT command).

6.63.7 /TEMPORARY

If specified the PROFILE changes given are only effective for the current NEWS session.

6.64 SET SEEN

By default, this command marks items within the currently selected newsgroup as ”seen”. This can be altered by specifying one of the following qualifiers.

6.64.1 /POINTER

This command marks all newsitems up to (and including) the ’current’ item as having been viewed by the user.

6.64.2 /BEFORE=date

This command marks all items in the current newsgroup which were postmarked by the local system before the date specified as having been viewed by the user.

The date format is the standard VMS date-time format; e.g. SKIP/BEFORE=22-FEB. ”TO-DAY” and ”YESTERDAY” are accepted as valid date strings.

6.64.3 /FOLLOWUP

All newsitems in the current newsgroup with a subject field the same as the current newsitem will be marked as read.

This command is used to skip entire conversation streams: all newsitems with related subject fields will be marked as having been read.

6.65 SHOW CLASSES

This command lists all user defined class names, and also lists the newsgroup names which have been assigned to each class. The currently selected class is marked with an asterisk.

6.66 SHOW CONFERENCE

Format:

    SHOW CONFERENCE newsgroup-name

Displays the status and configuration information for the specified newsgroup. This information includes the current item count, the default retention time settings, the markers placed within the newsgroup and the classes into which the newsgroup has been added, as well as the relevant NNTP server information if the newsgroup is configured as NNTP-served.

The newsgroup name be be abbreviated to a unique substring, and wildcards (’*’, ’%’) may be used to list the attributes of a number of newsgroups.

This command is equivalent to SHOW ENTRY.

6.67 SHOW ENTRY

Format:

    SHOW ENTRY newsgroup-name

Displays the status and configuration information for the specified newsgroup. This information includes the current item count, the default retention time settings, the markers placed within the newsgroup and the classes into which the newsgroup has been added, as well as the relevant NNTP server information if the newsgroup is configured as NNTP-served.

The newsgroup name be be abbreviated to a unique substring, and wildcards (’*’, ’%’) may be used to list the attributes of a number of newsgroups.

6.68 SHOW ITEM

Format:

    SHOW ITEM [item-num]

SHOW ITEM displays the retention settings for the current or specifies newsitem.

6.69 SHOW KILL

This command displays the current KILL filter set. Kill filters are applied to all unseen items, and if an item matches any of the kill filters the item will not be displayed when using the READ/NEW command.

6.70 SHOW MARKER

Format:

    SHOW MARKER [tag-value]

SHOW MARKER displays a listing of the set of marked newsitems. If no tag value is specified, then all mark tag values are displayed. If a tag is used then only those tag values which match the tag (using wildcard matching) are displayed. See Also MARK, UNMARK, CLEAR MARK

6.71 SHOW MEMBER

Format:

    SHOW MEMBER [username]

This command lists the attributes of a member of the currently selected newsgroup. If the newsgroup is a public access newsgroup then there is no associated membership list, and a status message is displayed.

The listing displays the username of the member, the user’s privileges with respect to this newsgroup and their mail addresses (if not that of their local username).

If /ALL is not specified, and no username is given, NEWS will prompt for a username.

6.71.1 /ALL

List the attributes of all members of a newsgroup.

6.71.2 /MODERATOR

List the attributes of all members who have moderator privilege with respect to this newsgroup.

6.72 SHOW MODERATOR

This command displays the mail address of the moderator of a moderated newsgroup. The newsgroup referred to by this command is the currently selected newsgroup.

The local NEWS manager may also nominate a set of local users who have moderator privileges for this newsgroup. Such privileges all the listed users the ability to alter the membership list of the newsgroup and a number of other newsgroup control functions. This command also lists the usernames of such local moderators and their mail addresses.

6.73 SHOW NEWSGROUP

Format:

    SHOW NEWSGROUP [newsgroup-name]
    SHOW NEWSGROUP *

SHOW NEWSGROUP displays the control attributes of the current or specified newsgroup. The control attributes include the retention settings, server settings and moderator address.

6.74 SHOW NOTE

Format:

    SHOW NOTE [item-number]

This command displays the header lines of the current (or nominated note if an item number is specified with the command. As well as the supplied header lines, the display also shows the local postmark time, and the local expiration date of the item.

6.75 SHOW PROFILE

This command displays your current NEWS profile settings. The display includes the default editor, the default class on startup, the NEWS personal name, default qualifiers for the PRINT command, OPEN default action and the startup directory setting. Note that PROFILE entries are at a higher precedence than the logical names which also define your NEWS profile.

See Also SET PROFILE

6.76 SHOW VERSION

This command displays the software version and creation date of the local NEWS software.

6.77 SKIM

SKIM is a NEWS MANAGEMENT command.

SKIM checks aspects of the news database for internal consistency, and also enforces news item and newsgroup expiry time limits.

By default the command will assume the qualifiers /ITEMS (check all news items), /NEWS-GROUPS (check all newsgroups), /SERVED (check all remote servers for consistency with the local database), /DIRECTORIES (check all directories in the database area), /HISTORY (skim the history file) and /FILES (check all item text files).

Any of these phases can be bypassed by using the /NOphase qualifier with the SKIM command.

Skim optionally accepts one parameter, a list of newsgroups to examine. If no parameter is given, then all newsgroups are checked. The parameter is a comma separated list of newsgroup names. Wildcard characters (’*’, ’?’) may be used to specify a set of newsgroups. e.g. ”SKIM aus.*,anu.*”

6.77.1 /DIRECTORIES

NEWS stores news items in a directory tree, where each directory matches a newsgroup: for example all items in the newsgroup ”aus.fred” are stored in the directory NEWS_DEVICE:[AUS.FRED]. When a newsgroup is deleted, the associated directory is not deleted immediately.

The SKIM/DIRECTORIES phase scans all directories in the NEWS area, and deletes all directories which do not match currently defined newsgroups.

6.77.2 /FILES

By default the action of the SKIM command includes SKIM/FILES. In this phase of the skim, SKIM examines every .ITM file (newsitem text file) in the news database, purging all back versions of any file, and deleting any .ITM files which have no corresponding entry in the item database.

This action can be bypassed using the /NOFILES qualifier on the SKIM command.

6.77.3 /HISTORY

Format:

    SKIM/HISTORY[=n]

NEWS maintains a history file of the previous n days, storing all newsitem message identifiers in this file. This file is used to correctly implement the ”cancel” and ”supercedes” news header control lines, and to detect loops on the NEWS topology.

This phase of SKIM removes all entries from the history file which are older than n days. The default retention if not specified is 45 days.

6.77.4 /ITEMS

By default the action of the SKIM command includes SKIM/ITEMS. In this phase of SKIM, the procedure checks all current newsitems held in the local database. If the item has been held for longer than the locally set retention period, then the item is deleted, or if the text file cannot be opened the item is deleted.

The items to be deleted can optionally be archived using the /ARCHIVE qualifier.

This action can be bypassed using the /NOITEMS qualifier with the SKIM command.

6.77.5 /NEWSGROUPS

By default, the action of the SKIM command includes SKIM/NEWSGROUPS. In this phase of SKIM, skim checks each newsgroup, updating the newsitem count for the newsgroup, and deletes any newsgroup which has had no new items posted in the retention period (see SET NEWSGROUP/HOLD=n for more information on newsgroup retention).

This action can be bypassed using the /NONEWSGROUPS qualifier on the SKIM command.

6.77.6 /SERVED

By default, the action of the SKIM command includes SKIM/SERVED. In this phase of skim, SKIM collects the item identifiers of all newsitems whose text is held on the remote NEWSSERVER node. SKIM then connects to the remote server and collects all item update information from the remote item database. Following this remote interrogation, the local item database is updated to match that of the remote node.

This action can be bypassed using the /NOSERVED qualifier on the SKIM command.

6.77.7 /ARCHIVE=filename

By default the value of this qualifier is /NOARCHIVE.

This qualifier can be used in conjunction with the /ITEMS qualifier to archive a copy of every item that is to be deleted from the News database. The items are stored using batch format (items are delimited by a line of the form ”#!rnews <size>. The items are appended to the archive file until the archive file reaches 250k bytes in size, when a new version of the archive file is opened.

6.78 SKIP

SKIP is used to mark individual newsitems or all newsitems within a newsgroup as having been read by the user without viewing the item texts.

If the screen is displaying the newsitem directory, then SKIP marks the ’current’ newsitem as read, and advances the newsitem pointer to the next newsitem.

If the screen is displaying the newsgroup directory screen, or the /NEWSGROUP qualifier is specified, then ALL newsitems within the ’current’ newsgroup are marked as having been read.

6.78.1 /ALL

This qualifier will mark ALL news items in ALL newsgroups as having been read by the user.

6.78.2 /POINTER

This command marks all newsitems up to (and including) the ’current’ item as having been viewed by the user. This qualifier may only refer to the current newsgroup.

6.78.3 /BEFORE=date

This command marks all items in the current newsgroup which were created on the local system before the date specified as having been viewed by the user.

This qualifier may also be used with the /ALL qualifier to mark all items in all newsgroups, created before the specified date, as having been viewed.

The date format is the standard VMS date-time format; e.g. SKIP/BEFORE=22-FEB. ”TO-DAY” and ”YESTERDAY” are accepted as valid date strings.

6.78.4 /FOLLOWUP

All newsitems in the current newsgroup with a subject field the same as the current newsitem will be marked as read.

This command is used to skip entire conversation streams: all newsitems with related subject fields will be marked as having been read.

6.78.5 /NEWSGROUP

This qualifier will mark ALL items in the current newsgroup as read. If the screen is displaying the newsgroup directory then /NEWSGROUP is the default action - if the screen is displaying the newsitem directory then /NONEWSGROUP is the default action, and only the current item will be marked

6.79 SPAWN

Format:

    SPAWN [command-string]

Spawn a DCL level subprocess from NEWS (using LIB$SPAWN). LOGOUT from the subprocess returns to NEWS. If a command string is specified, then this command is executed and then control is passed back to NEWS.

6.80 TOPIC

TOPIC displays the original topic of the current item. This command refers to conversation streams, where each posting in a conversation contain explicit references to the preceeding articles in the same conversation.

TOPIC attempts to follow this chain of reverse references back to the original posting at the head of these backward references, or if this posting is no longer available, locates the oldest item in the conversation and display it.

READ/TOPIC is a synonym for TOPIC.

6.80.1 /HEADER

Directs NEWS to display all header lines of the newsitem. By default only the subject line, sender address and date of the item are shown in the display.

6.80.2 /EDITOR

Use a callable editor in read-only mode to display the newsitem. By default this is the TPU editor, but this may be altered by the NEWS_VIEW logical name. See the chapter on customizing NEWS for details. For compatibility with previous versions of NEWS, /TPU is equivalent to /EDITOR.

6.80.3 /ROT13

Some newsitem are posted using rot13 encryption. This qualifier decrypts the newsitem for display.

6.81 UNMARK

Format:

    UNMARK [tag]

UNMARK clears the mark from the current news item. If a tag is specified with the command, then only those tag values will be cleared. If no tag is specified, then all tags will be removed from the item. The tag may be specified using wildcard characters. See Also MARK, CLEAR MARK, SHOW MARK

6.82 UNREAD

This command marks the current item as UNREAD - The item will be remarked again as unread, and will be displayed at a later stage by the READ/NEW command (if the item is in a registered newsgroup).

This command is useful to mark items for re-reading at a later stage.

6.83 UP

Format:

    UP [number]

In screen mode this moves the cursor up one line. The command has an optional numeric parameter, the number of lines to move up. This command is bound to the up arrow key on the terminal, and the command UP 18 is bound to the PREV SCREEN key. UP is also bound to Keypad 5 by default.

6.84 UPDATE

Format:

    UPDATE [newsgroup]

When NEWS is invoked it takes a snapshot of the local NEWS database, and presents this snapshot in a directory format to the user. If another user posts to NEWS, or the NEWS transport system reads in a news batch during your session your local copy of the database summary is not updated.

UPDATE clears the image section which maps the database and reads in a new copy of the database snapshot, including any changes made since your session was started.

If a newsgroup name is specified then the update function is performed on that newsgroup only.

6.85 VERSION

This command displays the software version and creation date of the local NEWS software.

6.86 WRITE

WRITE is synonymous to the POST command.

Format:

    WRITE [filename]

WRITE is used to post a new news item into NEWS. By default, the editor is invoked to form the new posting. The default editor is TPU, but this is user definable - see the chapter on customisation to specify a different editor to use.

NEWS will fill in all defaulting news headers, and will prompt for all header items which are non-defaulting (the subject line and the newsgroups to post the item into).

If the newsgroups for the posting are moderated newsgroups, the posting is automatically redirected as mail to the moderator address, and NOT posted directly to NEWS.

If a filename is specified, the editor is loaded with the text of the specified file, otherwise the editor is invoked to create a new file.

See Also ANSWER, FOLLOWUP, FORWARD, POST, REPLY

6.86.1 /EDIT

The default action is to call the editor to create the posting. If you specify a file as a parameter to the WRITE command, and also specify the qualifier /NOEDIT , no call is made to the editor.

6.86.2 /CONTROL

Format:

    WRITE/CONTROL="string"

/CONTROL is a NEWS MANAGER qualifier. The string specified as the value of this qualifier is inserted into the posting as a ”Control:” header line argument.

6.86.3 /HEADERS

Write will normally prompt for the subject of the posting, and a list of newsgroups to post the item into. All other header lines of the news item text will be filled with default values.

The /HEADERS qualifier will force prompting for optional components of the item header. These fields are:

Summary: A one line summary of the contents of the posting.

Reply-To: The (internet) mail address for REPLYing to the item - used if replies are to be sent to a mail address OTHER than yourself (as the originator of the item.

Distribution: A (comma separated) list of keywords designed to limit the distribution of a news item to a subset of the network. These keywords are site specific, but commonly include the keyword ”local” (do not forward to the net!).

Followup-To: The list of newsgroups where followup postings are to be sent. This is used to move a discussion to another newsgroup.

Keywords: A few relevant, pithy keywords which should indicate to other readers if the item is of interest.

Expires: A date to indicate the end of the useful life of the information in the message. NEWS will reject expirations greater than 12 months as being too outlandish.

6.86.4 /NEWSGROUPS

Format:

 WRITE/NEWSGROUPS=newsgroup-list

This qualifier specifies the list of newsgroups to receive the posting. If this qualifier value is not specified, the user is prompted for the newsgroups. The list is a comma separated list of newsgroup names.

6.86.5 /SUBJECT=subj

Format:

 WRITE/SUBJECT="text string"

Specifies the subject of the message for the heading. If not specified, the user is prompted for the /SUBJECT value.

6.86.6 /SELF

If specified a copy of your posting is mailed to yourself via VMS MAIL.

Moderated Newsgroups

This chapter describes how moderated newsgroups are configured and maintained.

A moderated newsgroup is one where posting to the newsgroup is barred to all users other than the nominated moderator. Only the moderator may post new items into the newsgroup, and only the original sender, the moderator and the local system manager may post a cancel referring to an item in a moderated newsgroup.

Where users attempt to post to a moderated newsgroup, the posting is automatically re-directed to the newsgroup moderator address as mail addressed to the moderator’s address.

7.1 Management of Moderated Newsgroups

The local news manager is responsible for defining moderated newsgroups, and also specifying the address of the moderator. The management commands are the SET NEWSGROUP and the CREATE commands, specifying the qualifier /MODERATOR, with the moderator’s address as the value of this qualifier.

Some network-wide newsgroups are also moderated newsgroups. This is managed by the ”checkgroups” control message, which specifies which newsgroups are moderated (but note that this message will not specify the actual moderator address).

The local specification of moderated newsgroups is held in two areas of the NEWS database: each newsgroup descriptor contains a flag indicating whether the newsgroup is moderated or not, and a second file (the MAILPATHS file) contains the addresses of all newsgroup moderators.

The format of the ”mailpaths” file, containing the moderated newsgroup names, and the internet address of the moderator. However this file does not contain a comprehensive list of newsgroup moderators for the network-defined newsgroups: when examining the support structures for the USENET news network, one of the relevant aspects of the network is the presence on the net of a number of ”backbone” sites. These backbone sites contain an up-to-date list of the moderators of all network newsgroups, and are also configured to forward all moderated newsgroup postings on to the correct moderator. The scheme used in addressing the moderator via a backbone site is one where a posting to the moderated newsgroup ”comp.animals” (say) is directed to the address ”comp-animals@<backbone_node>”. The reserved newsgroup name ”backbone” matches all newsgroups, and the address may contain the string ”%s”, which requires the actual newsgroup name to be substituted, with ’_’ replacing all ’.’ characters in the newsgroup name. The scan of the file is first match found.

Once the ”mailpaths” file is initialized with the backbone entry, all further modifications to the file can be performed within NEWS, using the SET NEWSGROUP command.

7.2 POST

The POST command will normally re-direct all postings to a moderated newsgroup to the moderator (as mail).

However if the user’s internet address (built by using the logical name translation of NEWS_ADDRESS in the format ”<username>@<news_address>”) matches the moderator address, the posting is allowed, and the posted item is marked as ”Approved” so that the item will be distributed correctly through the network.

7.3 ADD FILE/MODERATOR

This is a variant of the ADD FILE command, and is the method whereby a newsgroup moderator may add new items into the newsgroup. This differs from the POST command in so far as the ”From:” address is preserved as the original submitter when using the ADD command, whereas POST assumes the posting was originated by the moderator.

If the original submitter posted to a number of moderated newsgroups, then the procedure will request confirmation to forward the item to the next moderator in turn, or if the item is already approved by all other moderators the posting will then be placed in all the moderated newsgroups specified by the original submitter.

Other optional qualifiers to the ADD command are documented in the command reference section of this manual, and are available with the /MODERATOR qualifier.

7.4 CANCEL

The CANCEL command is also altered by a moderated newsgroup, in so far as the moderator of the newsgroup may issue the CANCEL command (as well as the original submitter).

(

NEWS MANAGEMENT)

NEWS management is a two level hierarchy: management of individual newsgroups can be set as the responsibility of various local newsgroup managers (or local moderators), while overall management of NEWS is the responsibility of a news manager.

8.1 NEWSGROUP MANAGEMENT

A user is nominated as a local newsgroup manager using the ADD MEMBER command, granting the specified user ”MODERATE” privilege. The user can then access newsgroup management commands by first enabling the privilege using the SET MODERATOR command.

A newsgroup moderator can modify the access list of a newsgroup (ADD MEMBER, DELETE MEMBER, MODIFY MEMBER) and can also cancel any local postings made into the managed newsgroup (CANCEL). The management domain of a local newsgroup manager does not extend outside the local system, and all operations accessible to the local newsgroup manager only affect the local node.

8.2 (

NEWS MANAGEMENT)

NEWS managers are identified to NEWS by holding a site-specific identifier. A user with this identifier can access all newsgroups regardless of access restrictions, and can issue all management commands, including local newsgroup management commands. All management commands check to determine whether the user is a NEWS manager before executing the command.

The news manager normally controls the execution of the batch jobs which update and skim off the local NEWS database (ADD and SKIM commands), and manage attributes of the local database (SET commands). All these commands are documented in the command reference section of this document, with a note to the effect that the command is a NEWS MANAGEMENT command.

The responsibility of the local NEWS manager also includes that of the maintenance of the various site-specific control files. The files and their function are documented in the chapters describing the NEWS distribution set and NEWS installation.

NEWS Distribution

The software is distributed in source and executable formats. The following is an annotated directory listing of the NEWS V5.7 distribution file sets:

9.1 SRC distribution

The SRC kit contains the sources to the NEWS package. The files in that kit are as follows. VMS system definition files:

SJCDEF.H – SYS$SNDJBC item list definition module. SMGDEFEXTRA.H – Sundry SMG definitions missing in VMS C supplied definition files. SORTDEF.H – Definitions for the callable sort (SOR$) routines. V5MAILDEF.H – Definition of callable MAIL item codes and description of the VMS V5.0

Mail profile file and the MAIL.MAI file.

NEWS utility definition files. These are the common include files added to the start of all NEWS source code modules. Together they implement a consistent compile-time environment for all modules that are used to build NEWS.

NEWSDEFINE.H – Common constant and macro definitions. NEWSEXTERN.H – Common external variable definition for source modules. NEWSINCLUDE.H – Common include file set for all source modules. NEWSSITE.H – NEWS compile-time constants - may be altered to suit local requirements. NEWSVARIABLES.H – NEWS global variables definition module. NNTP.H – NNTP definitions The following set are the C sources to the main NEWS program. The modules are split on the basis of the commands that each module implements.

NEWS.C – NEWS entry point and outer control loop. NEWSADD.C – Add command routines. NEWSCREATE.C – Create command routines. NEWSDELETE.C – Delete command routines. NEWSDIR.C – Directory command routines. NEWSDISPLAY.C – Display item text on screen. NEWSEXTRACT.C – Extract/Print command routines. NEWSFORWARD.C – Forward/Reply/Mail command routines. NEWSHELP.C – Help command routines. NEWSMODE.C – Set terminal mode routines. NEWSPOST.C – Post/Followup command routines. NEWSREAD.C – Read command routines. NEWSREGISTER.C – Register/Deregister command routines. NEWSREMCLIENT.C – Utility routines to access remote NNTP Server. NEWSSEARCH.C – Search command routines. NEWSSELECT.C – Select command routines. NEWSSETSHOW.C – Set and Show command routines. NEWSSITE.C – Loading site specific logical name values, and routines for network distribution

of news to adjacent sites. NEWSSKIM.C – Skim command routines. NEWSSKIP.C – SKIP command routines. NEWSUTILITY.C – Various utility routines for NEWS. ADD_TRANSFORM.C – Translation from internet address format to one acceptable to VMS

MAIL.

The NEWS command language definition file:

NEWSCMD.CLD – NEWS command language definition file (NEWS uses the CLI parser command parsing).

The C source modules for the NNTP implementation:

NNTP_CLIENT.C — Interactive access to remote NNTP server. NNTP_DEC.C – DECNET driver for NNTP server. NNTP_DECM.C – DECNET driver for NNTP server, with support for multiple DECnet con-

nections to a single server process. NNTP_SERVER.C – Source for NNTP NEWS Server object. NNTP_TCPCMU.C – TCP driver for NNTP server, for the CMU/Tek implementation of

TCP/IP for VMS. NNTP_TCPWIN.C – TCP driver for NNTP server, for the WIN implementation of TCP/IP

for VMS. NNTP_TTY.C – TTY (interactive access) driver for NNTP server. NNTP_FEED.C – NNTP processor for NNTP ihave-based NEWS distribution.

NEWS On-line help library source:

NEWS.HLP – Help library source

Compilation and link files:

NEWSBUILD.COM – DCL command procedure to perform a makefile of NEWS OPTIONS_C_LINK.OPT – VMS LINK options commands

9.2 OBJ distribution

The OBJ distribution contains the object library modules and the ”dummy” object files used in the BUILDNEWS command procedure. Note that many of the OBJ files are actually text files - these files are used only to determine the date and time of the last change to the corresponding module within the object library files, and hence do not need to contain the actual compiled object code. To save space the BUILDNEWS procedure replaces the object file with an object marker file following the change to the object library file. The .OBJ files are not listed here. The two object library files are:

NEWS_LIBRARY.OLB – object modules compiled with OPT and NODEBUG NEWS_LIBRARYD.OLB – object modules compiled with NOOPT and DEBUG

9.3 DIST distribution

This distribution contains the executables and the associated DCL command procedures, control template files, and with release information.

Release information:

AAAREADME.TXT – distribution documentation CHANGES.TXT – brief description of changes to NEWS from previous versions BUGS.TXT – description of known bugs NEWS.DOC – This file

Executable files:

NEWS.EXE – NEWS utility, compiled under VMS V4.7 NNTP_DEC.EXE – NNTP server, DECnet transport NNTP_DECM.EXE – NNTP server, DECnet transport, multithreaded server NNTP_TCPCMU.EXE – NNTP server, CMU/Tek TCP/IP transport NNTP_TTY.EXE – NNTP server, interactive interface NNTP_CLIENT.EXE – NNTP interactive client (test utility) NNTP_FEED.EXE – NEWS network transport utility, NNTP protocols

On-line HELP file

NEWS.HLB – On-line help library

All the files in the next list are site-specific template files, which MUST be edited before installing on the local system. Each of these files contains detailed descriptions of how the file is to be edited, and what aspects of NEWS are controlled by the contents of the files.

MAILPATHS – Template of the control file used to generate the internet addresses of moder-

ators of newsgroups. NEWS.DISTRIBUTION – Template of news distribution control file – used to determine the

network routing of newsitems between adjacent news network neighbours. NEWS.SYS – Template of the control file used to determine which newsgroups and distribution

keywords are accepted by the local NEWS site, and which newsgroups and distribution keywords are accepted by the local node’s adjacent NEWS network neighbours.

SETUP_LOG_EXAMPLE.COM – Template file for NEWS logical name definitions. NEWSSKIM.COM – Template automatic news management routine (designed for periodic

execution in a batch stream). NEWSADD.COM – Component of NEWSSKIM which handles addition of network batches

and distribution to downstream sites.

NEWSDAILY.COM – Component of NEWSSKIM which performs expiration of items and

internal consistency checks on the news database. NEWSPACK.COM – Component of NEWSSKIM which calls CONVERT to compress the RMS

indexed files. NEWSEDIT_EXAMPLE.COM – Template file for invoking any editor from NEWS. NEWS_ADDRESS.CNF – Template file for translation rules to map internet-style addresses

to a form acceptable to the local VMS MAIL system. NEWS_POST.CC – Template file for the automatic generation of carbon copies of all local

postings. NEWS_POST.DEFAULTS – Default header definition file for local postings NNTP_FEED.COM – Template for passing NEWS distributions to adjacent nodes using the

NNTP protocols.

9.4 DOC distribution

NEWS.LATEX – LaTeX source of this document. NEWS.MSS – Scribe source of this document. FILES.MSS – Component of NEWS.MSS - description of distribution files and installation

procedures. FILES.TEX – Component of NEWS.LATEX - description of distribution files and installation

procedures MGR.MSS – Component of NEWS.MSS - description of management commands. MGR.TEX – Component of NEWS.LATEX - description of management commands. USER.MSS – Component of NEWS.MSS - description of user commands. USER.TEX – Component of NEWS.LATEX - description of user commands. RFC977.DOC – News Network Transfer Protocol standard. NEWS-STANDARD.DOC – Usenet news standard (pre-RFC1036). NEWS-STANDARD.ANNOTATED – Standard, annotated with capabilities and exceptions

related to this version of VMS NEWS. RFC1036.DOC – USENET News standard.

Installing NEWS

The following section describe the necessary steps to install NEWS on a system. Note that where possible these steps should be performed in the order as given in this text.

NOTE: NEWS is shipped with object and executable files included, but these files were built using VMS V4.7. If you are running a different version of VMS then minimally you will need to re-link the NEWS object files

10.1 Installing NEWS files

1. Load *.C, *.H, *.HLP, *.CLD, *.OPT and NEWSBUILD.COM into the directory device:[dir.news.news_src]. This set of files corresponds to the SRC backup saveset of the distribution. The latter two directory names, NEWS and NEWS_SRC are critical when using the supplied DCL news build routines.

2. Create the directory device:[dir.news.news_build]. Load the contents of the OBJ backup saveset into this directory.

3. Create the directory device:[dir.news.news_dist]. Load the EXE backup saveset into this directory.

4. Create the directory device:[dir.news.news_build]. Load the DOC backup saveset into this directory.

10.2 Compiling and Linking NEWS

One command procedure performs both the compile and link phases of NEWS. The procedure works on the revision date-time of the file to determine whether any files require re-compilation before linking.

Execute the newsbuild command procedure:

    $ SET def dev:[dir.NEWS.NEWS_BUILD]
    $ @[-.NEWS_SRC]NEWSBUILD

(The SET DEFAULT COMMAND is critical to the build procedure)

This procedure creates a new set of .EXE files in the NEWS_DIST directory. which are used in the following installation steps.

10.3 NEWS Accounts

Add an account for a NEWS manager (NEWSMGR if you would like a username suggestion). This account runs the housekeeping batch jobs (such as newsskim if you are using it) and also runs the network feeds to adjacent NEWS sites, and therefore should have proxy entries to adjacent DECnet nodes’ NEWS manager accounts if more than one node will be running NEWS on the DECnet.

The account normally requires SYSPRV to execute some of the functions associated with the ’housekeeping’ batch jobs. The account may also need a reasonable disk quota if you are anticipating a large amount of network news flowing through your system, and you make this account own the news posting directories.

10.4 Logical Names

It is intended that much of the local tailoring of NEWS is performed by logical name translations. For this reason the logical name structure is important when installing NEWS - please read this section carefully.

NOTE: It is important that these names be defined before running news for the first time.

The distribution set includes the following file: setup_log_example.com

$ ! SETUP_LOG_EXAMPLE.COM
$ !
$ ! Logical name definition template - this will need to be altered to
$ ! match local site requirements
$ !
$ ! Author:
$ !     G Huston
$ !       Computer Services Centre
$ !       Australian National University
$ !
$ ! Version:
$ !     V1.0    17-Jul-1986
$ !     V2.0    16-Jul-1987
$ !     V4.3    12-Jan-1987     GIH
$ !     V5.2    26-Apr-1988     GIH
$ !     V5.6     3-Nov-1988     GIH
$ !     V5.7     3-Jan-1989     GIH
$ !      -  Added entries for rnews and newsmail mail addresses
$ !

$ !**********************************************************************
$!
$! root of news tree
$!
$ DEFINE/SYSTEM/EXEC NEWS_ROOT ANUDISK:[NEWS]
$ DEFINE/SYSTEM/TRANS=(conc)/EXEC NEWS_DEVICE CSC3$DUB0:[NEWS.]
$!
$! News manager accounts must own this identifier
$!
$ DEFINE/SYSTEM/EXEC NEWS_MGR_ID NEWSMANAGER
$!
$! the node address is the internet address of the node
$!
$ DEFINE/SYSTEM/EXEC NEWS_ADDRESS "CSC3.ANU.OZ"
$!
$! the node name is either the decnet node name, or cluster decnet alias
$!
$ DEFINE/SYSTEM/EXEC NEWS_NODE CSC3
$!
$! The "work directory" for news - normally sys$login of the
$! news manager account
$!
$ DEFINE/SYSTEM/EXEC NEWS_MANAGER ANUDISK:[NEWS_MGR]
$ DEFINE/SYSTEM/TRANS=(conc)/EXEC NEWS_MANAGER_DEV CSC3$DUB0:[NEWS_MGR.]
$!
$! Organisation definition
$!
$ DEFINE NEWS_ORGANISATION "CSC, Australian National University"
$!
$! Mail addresses
$!
$ DEFINE/SYSTEM NEWSMAIL NEWSMGR
$ DEFINE/SYSTEM RNEWS NEWSMGR

Description of names:

NEWS_ROOT, NEWS_DEVICE Both names point to the news files root directory. Both must be defined in exec mode. The news data files are placed in a tree descending from this root - i.e. newsgroup ‘aus.wanted’ has text files residing in the area news_device:[aus.wanted]. The Index files and Help file are placed in the news_root directory.

NEWS_MGR_ID Translates to the identifier name which must be granted to any users who require the management functions of NEWS (delete newsgroup, create newsgroup, etc). The name must be defined in exec mode to stop unauthorized users tinkering with NEWS.

NEWS_ADDRESS Translates to the internet address of the host node. This is typically of the form node.site.domain.

NEWS_NODE Translates to the local node name. This is made a logical name to allow a cluster to operate as a single logical node, using a cluster alias. All nodes in a cluster should use the same value of NEWS_NODE.

NEWS_ORGANIZATION Translates to the name of the organisation. This is to help in identifying the person posting the message to other users on the net. Users may use a process level logical name to overide this system value.

NEWS_MANAGER, NEWS_MANAGER_DEV This is the work directory of the NEWSMGR account. All temporary files used in news maintenance are placed here. NEWS_MANAGER_DEV is the equivalent area, defined as a concealed device.

NEWSMAIL, RNEWS These are mail address aliases to allow this node to communicate to adjacent NEWS nodes. They should be defined as the username of the NEWS manager account. An alternative is to use SYS$EXAMPLES:MAILUAF.COM, add an entry for the addresses ”NEWSMAIL” and ”RNEWS” with a forwarding address to the NEWS manager account (NEWSMGR).

10.5 Global DCL Symbols

In the system login command add a symbol for NEWS:

$ NEWS :== $ location:NEWS

10.6 Install NEWS.EXE

Copy NEWS.EXE to the location as specified by the DCL symbol for NEWS. Then install the image with SYSPRV.

$ INSTALL
INSTALL> ADD location:NEWS.EXE/OPEN/HEAD/SHARE/PRIV=SYSPRV

SYSPRV is required to ensure that users who post news items end up NOT owning the actual news text file, and also is required to successfully initialise the indexed files on the first-time execution of the program.

10.7 Initial Environment

• Using AUTHORIZE, create the NEWSMANAGER identifier (the value of the logical name NEWS_MGR_ID), and grant the identifier to the user(s) who will be responsible for the management of NEWS. the NEWS manager account created in the second step must be granted this identifier.

• Create the directories that have been defined by the logical names NEWS_MANAGER and NEWS_ROOT. [Optionally] Add a propagation ACL to the NEWS_ROOT directory to allow those users whom you wish to use news to access the item files. Ensure that the directories are group and world accessible.

• Using SYS$EXAMPLES:MAILUAF.COM, add an entry for the address ”USENET”, with a forwarding address to the news manager on your host. See also the note on the addresses ”RNEWS” and ”NEWSMAIL” above.

• Copy NEWS.HLB to NEWS_ROOT: and allow world access to this file.

10.8 NEWS Control Files

There are six control files – these files should be read, and then edited to match the local site requirements. Each file contains comments as to the function and format of the file.

NEWS.SYS Read (and then EDIT) file NEWS.SYS. This file determines which newsgroups you will accept, and which newsgroups your network neighbours will accept. The edited version of NEWS.SYS should be copied to NEWS_MANAGER:NEWS.SYS. Include the entries ”local”, ”junk”, ”control” and ”to.<node>” (where <node> is the value of NEWS_NODE) in the local node’s filter set.

NEWS.DISTRIBUTION Read (and then EDIT) the file NEWS.DISTRIBUTION. This file determines how news is distributed through your node. If you are an ’end-node’ for NEWS, then this file should be empty, otherwise the file must be defined. The edited version of NEWS.DISTRIBUTION should be copied to NEWS_MANAGER:NEWS.DISTRIBUTION.

MAILPATHS Read (and then EDIT the file MAILPATHS. This file defines the internet addresses of ALL moderators for moderated newsgroups. The edited version of MAILPATHS should be copied to NEWS_MANAGER:MAILPATHS.

NEWSSKIM.COM Read NEWSSKIM.COM modify it to suit your requirements. See comments in the code. Then copy the edited file to NEWS_ROOT. This file is responsible for the collection and distribution of news items to remote sites.

NEWS_ADDRESS.CNF Read this file, and also read the file ADD_TRANSFORM.C. If you are using a local PMDF mail handler, then the default action may be suitable. Otherwise you will need to edit th the .cnf file, and copy the file to news_manager:news_address.cnf.

NEWS_POST.DEFAULTS Read this file and decide if it is relevant. If so then edit the file and place a copy in news_manager:news_post.defaults.

NEWS_POST.CC Read this file, and decide if it is relevant. Where unmoderated newsgroups are fed from a mailing list, then postings into the newsgroup may not be sent back to the central mailing list address. You can place entries into this file to force cross mailings of local postings back to the mailing list address. If necessary, edit this file and copy it to news_manager:news_post.cc. If not required then do not copy the file to the news_manager area.

10.9 First Time Execution of NEWS

Once ALL the above steps have been completed, login to an account which has SYSPRV and has the NEWS_MGR_ID identifier, and run NEWS.EXE.

The first time execution will create the files: NEWS_ROOT:NEWSGROUPS.V50 and NEWS_ROOT:NEWSITEMS.V50, and also create the file NEWS_MANAGER:NEWS.LICENSE.

Please read the NEWS LICENSE file, edit it as instructed, and return it to the address shown in the file.

This almost completes the installation of NEWS - the remaining task is the creation of the necessary newsgroups. Use the CREATE/LOCAL command and create the newsgroups ”control” and ”junk”. If you are on a news network, obtain a copy of the most recent ”checkgroups” posting. Examine the newsgroup names and add the root name (e.g. ”net” from ”net.wombats”) of the relevant newsgroups to your local acceptance filter in NEWS.SYS. Once NEWS.SYS has been edited, and pass the checkgroups file to NEWS (using the ADD command). This will mail a command procedure to the MAIL address ”USENET” which will setup NEWS with the current network newsgroup set. Then add any local newsgroups to this set using the CREATE/LOCAL command.

Upgrading from Previous NEWS Versions

If you are running NEWS Versions 4.0, 4.1, 4.2 or 4.3 then the upgrade is a two step process: • Using a SYSPRV account run the program NEWS V43 V44.EXE.

• This will create the file NEWS_ROOT:NEWS_V4.GROUPS. Delete the file NEWS_ROOT:NEWS.GROUPS.

• Edit the file NEWS_MANAGER:MAILPATHS to suit local requirements

• Follow the V4.4 to V5.0 instructions. If you are running NEWS V4.4 then:

• Using a SYSPRV account run the program NEWS_V44_V50.EXE.

• This will create the file NEWS_ROOT:NEWSGROUPS.V50 and NEWS_ROOT:NEWSITEMS.V50.

Delete the files NEWS_ROOT:NEWS_V4.GROUPS and NEWS_ROOT:NEWS.ITEMS. If you are running NEWS V5.0 or NEWS V.51 then:

• Execute the command procedure NEWS_V51_V52.COM, then Purge old versions of newsitems.v50

in news_root:. If you are running NEWS V5.4 or NEWS V5.6 then:

• Delete the file NEWS_ROOT:HISTORY.DAT

Installing NNTP NEWS Servers

There are three node types with respect ot NEWS servers: Server nodes, where the node acts as a server for remote clients, Client-only nodes, which use one or more server nodes, and Server/Client nodes, where some newsgroups are served from remote servers, and this node acts as a server for remote client nodes.

If you wish to install the NNTP NEWS Servers, then follow the steps as shown for the node types. You can check the operation of the server by running the NNTP_CLIENT program (this program implements an interactive interface to these NEWS servers). Running NNTP_CLIENT and specifying the remote node name and transport protocol should initiate an interactive session with the remote NNTP server. The welcome message should appear, and you should then obtain a reasonable response to the ”HELP” command. The ”QUIT” command will exit from the remote server.

Once NNTP is installed on the server and client newsgroups can then be set as NNTP-served newsgroups (see the documentation on the SET command for details).

12.1 DECNET Server and Client Nodes

If you are installing a SERVER or a mixed SERVER/CLIENT the installation steps required for these node types are:

• Copy NNTP_DEC.EXE to a location accessible to the DECNET account.

• Use the INSTALL utility to install NNTP_DEC.EXE with SYSPRV.

• In the DECNET account create the file NNTP.COM with the following contents:

       $ ! NNTP.COM - executed in response to an incoming connect
       $ ! to "TASK=NNTP"
       $ RUN dev:[dir]NNTP_DEC

There is a variant of the NNTP_DEC.EXE program, called NNTP_DECM.EXE. This program is configured to accept multiple incoming DECNET connections to the NNTP object, and service all connections from a single NNTP process. The installation instructions are the same as the NNTP_DEC instructions, except that the executable referred to in the IN-STALL and NNTP.COM is NNTP_DECM.EXE. Also to create a permanent DECnet object, NNTP_DECM must be installed with SYSPRV and SYSNAM. (NOTE: I have not exhaustively tested NNTP_DECM, and there may well still be bugs in this code under certain conditions. I would be grateful for any comments on problems encountered with this server.)

For a CLIENT there is no special installation required (however if the server is running an earlier version of NEWS then the notes in the following section do apply).

12.1.1 Compatibility with Previous NEWS Versions

There are two major differences between earlier versions (V5.4 or earlier) and this version of NEWS: The server executable has been renamed from NEWSSERVER.EXE to the file NNTP_DEC.EXE, and the task connection string has been changed from ”NEWS=” to ”TASK=NNTP”. The first change was a prompted by a desire for a better name for the NNTP server image, and the second by a desire to achieve compatibility with the existing implementation of NNTP under Ultrix, where NNTP is mapped into a DECnet object ”TASK=NNTP”.

If you are upgrading ALL nodes in your network then the object name change presents few problems, as all nodes are changed to refer to the ”NNTP=” connection string.

If you are upgrading only the SERVER, and clients still exist which wish to connect to the task ”NEWS=”, then you’ll need to map NNTP_DEC.EXE to this task string. This is done by inserting the following entry into the volatile and permanent NCP databases:

  $ RUN SYS$SYSTEM:NCP
  NCP> SET OBJECT NEWS NUMBER nnn FILE dev:[dir]NNTP_DEC.EXE
  NCP> DEFINE OBJECT NEWS NUMBER nnn FILE dev:[dir]NNTP_DEC.EXE

Where the object number matches that chosen for the previous version of the NEWS object. This also applies when using NNTP_DECM on the server - the .EXE name is changed in the above recipe.

If you are upgrading the CLIENT, and the server only accepts connections to ”NEWS=”, then the steps are:

  $ ! node in the following line refers to the node name of the NNTP
  $ !server on DECnet: e.g.
  $ ! define/system news_csc_task "NEWS="
  $ !
  $ DEFINE/SYSTEM NEWS_node_TASK "NEWS="
  $ RUN SYS$SYSTEM:NCP
  NCP> SET OBJECT NEWS NUMBER nnn FILE dev:[dir]NNTP_DEC.EXE
  NCP> DEFINE OBJECT NEWS NUMBER nnn FILE dev:[dir]NNTP_DEC.EXE

This logical name: NEWS_node_TASK is used to replace the default task specification of ”TASK=NNTP” with the supplied task string when connecting to that particular node.

12.2 CMU TCP Server and Server/Client Nodes

The installation steps required for these node types are:

• Copy TCPSERVER.EXE to a location accessible to the SYSTEM account. (this example uses sys$system)

• In sys$manager:internet.config insert the following line:

; Define the well-known ports we support and their servers
; WKS:port:procname:imagename:status:privs:priority:queuelimit
;
WKS:119:NEWSRV:TCP$NEWSRV:NETWRK:NETMBX,TMPMBX,SYSPRV,PHY_IO:4:5

• In sys$manager:ip_startup.com insert the following line:

$ define /system tcp$newsrv sys$system:tcpserver.exe

12.3 WIN/TCP Server and Server/Client Nodes

NO DETAILS HERE AS I DO NOT RUN WIN/TCP - CAN SOMEONE PROVIDE THEM PLEASE?

NEWS Feed to a Unix Site

This section describes some hints as to how to feed new postings from a VMS NEWS site to an adjacent Unix system. There are two basic approaches: Using NNTP and TCP/IP on VMS, or using a VMS/Unix mailer.

13.1 NNTP-based NEWS feeds

The NNTP approach uses the ihave/sendme protocol described in the NNTP standard. The following changes have to be applied:

Modify NEWS.SYS on the VMS system:

unixnode:world,inet,arp,anu,canb,comp,sci,rec,news,\
misc,soc,talk,aus,alt:N:\
NEWS_MANAGER_DEV/[IHAVE_UNIXNODE]COLLECT.IDS

The relevant entry now specifies that the VMS system will append the message identifiers to the specified file, rather than placing the entire message text into the posting area (the ’N’ flag in the SYS entry specifies message identifier transmission).

In the Skim command procedure on VMS, the following commands need to be inserted:

$! Copy news items to other sites
$distribute:
$ write sys$output "Copy items to other sites"
$ set on
$ NNTPFEED :== $ dev:[dir]NNTPFEED
$ search_1:
$ nodedir = f$search("news_manager:post_*.dir",2)
$ if nodedir .eqs. "" then goto abort
$ node = f$parse(nodedir,,,"NAME","SYNTAX_ONLY") - "POST_"
$ senddir = "news_manager_dev:[post_" + node + "];"
$ savedfiles = senddir - ";" + "*.*;*"

$ node = node + "::;"
$search_2:
$ file = f$search(savedfiles,3)
$ if file .eqs. "" then goto search_1
$ on error then goto no_copy_2
$! Detect you are transferring a batch to UNIXNODE:
$!        $ NNTPFEED unixnode decnet 'file ! for DECNET
$! - or-  $ NNTPFEED unixnode tcp 'file    ! for TCP
$ if node .eqs. "UNIXNODE::;" -
 then $ NNTPFEED unixnode tcp 'file
$ if node .nes. "UNIXNODE::;" -
 then $ copy/log 'file 'node
$ delete 'file
$ no_copy_2:
$ goto search_2

If the VMS system does not feed to additional systems then no further changes are necessary (i.e. the VMS system is a logical end-node in the news network topology). However if the VMS system does relay network batches (i.e. a routing node with respect to the NEWS network), then you will need to modify NEWS.DISTRIBUTION to ensure that news batches are forwarded correctly through this VMS node to all adjacent nodes.

13.2 Mail-based NEWS feeds

The second approach is based on the existance on a mail link between the two systems. On the VMS system you will need to define two mail destinations: ”newsmail” and ”rnews”. These mail destinations are used by the Unix system as the default mail destination of forwarded news batches. If the news manager username is ”NEWSMGR” then the following logical names need to be defined on the VMS system:

$ DEFINE/SYSTEM NEWSMAIL NEWSMGR
$ DEFINE/SYSTEM RNEWS NEWSMGR

The VMS NEWS.SYS file requires the entry:

unixnode:world,inet,arp,anu,canb,comp,sci,rec,news,\
misc,soc,talk,aus,alt:B:\
NEWS_MANAGER_DEV/[POST_UNIXNODE]NEWS.BATCH^BN

The ’B’ flag indicates the news item texts are to be stored in the post area, and the ’BN’ transmission flags indicate that the the mail format uses leading ’N’ characters in the mail body to delineate the news batch from the mail headers.

In the Skim command procedure on VMS, the following commands need to be inserted:

$! Copy news items to other sites

$distribute:
$ write sys$output "Copy items to other sites"
$ set on
$ search_1:
$ nodedir = f$search("news_manager:post_*.dir",2)
$ if nodedir .eqs. "" then goto abort
$ node = f$parse(nodedir,,,"NAME","SYNTAX_ONLY") - "POST_"
$ senddir = "news_manager_dev:[post_" + node + "];"
$ savedfiles = senddir - ";" + "*.*;*"
$ node = node + "::;"
$search_2:
$ file = f$search(savedfiles,3)
$ if file .eqs. "" then goto search_1
$ on error then goto no_copy_2
$ if node .eqs. "UNIXNODE::;" -
 then $ mail 'file' in%"""rnews@unixnode"""
$ if node .nes. "UNIXNODE::;" -
 then $ copy/log 'file 'node
$ delete 'file
$ no_copy_2:
$ goto search_2

On the Unix system an entry needs to be inserted in the aliases file:

rnews: "|/usr/lib/news/uurec"

Variants of these approaches are possible, such a mailing the ’ihave’ list of message-ids to the unix site, and then having the Unix site respond with ’sendme’ control messages back to the VMS system using the mailer as the message transport. The comments in the distributed version of NEWS.SYS contain additional information on how to set up such distribution mechanisms.